### v1

# Constructing embedded surfaces for cellular embeddings of leveled spatial graphs

Abstract
Finding a closed orientable surface $\mathcal{S}$ embedded in $\mathbb{R}^{3}$ where a given spatial graph $\mathcal{G}\subset\mathbb{R}^{3}$ cellular embeds is in general not possible. We therefore restrict our interest to the special class of spatial graphs that are leveled. We show that for leveled spatial graphs with a small number of levels, a surface $\mathcal{S}$ can always be found. The argument is based on the idea of decomposing $\mathcal{G}$ into subgraphs that can be placed on a sphere and on handles that are attached to the sphere, together forming an embedding of $\mathcal{G}$ in $\mathcal{S}$. We generalize the procedure to an algorithm that, if successful, constructs $\mathcal{S}$ for leveled spatial graphs with any number of levels. We conjecture that all connected leveled embeddings can be cellular embedded with the presented algorithm.
1 Introduction
Embeddings of graphs on surfaces are studied as central objects in a wide variety of fields. They are the main object of study in topological graph theory, where existence and properties of embeddings of given abstract graphs in abstract surfaces are considered, as well as in geometric graph theory, where the surfaces are endowed with a metric and the graphs are usually assumed to have straight-line edges. In knot theory, graph embeddings are used to produce invariants for knots, such as the Heegaard genus of a knot, which is the smallest genus of a handlebody the knot can be embedded in. In synthetic chemistry, graph embeddings are frequently used to describe the bond network of molecules.
Cellular embeddings, as a natural generalization of planar graphs, have in particular been studied in topological graph theory and combinatorics. They provide a combinatorial description of graph embeddings in surfaces [9], and they are at the core of long-standing open conjectures such as the circular (or strong) embedding conjecture [6].
It is possible to cellular embed any abstract graph in a closed orientable surface. For a given spatial graph, it is still always possible to find a closed orientable surface it embeds in by considering its regular neighborhood and pushing the spatial graph on the obtained surface. But in contrast to abstract graphs, it is not possible to find a cellular embedding for every spatial graph. For example, for nontrivial knots, there exist no such surface where the knot cellular embeds. This motivates the question for which spatial graphs a closed orientable surface embedded in $\mathbb{R}^{3}$ can be found, such that the graph cellular embeds in it, and how the surfaces can be constructed.
Given that graphs on surfaces have mainly be studies in topological graph theory that, in contrast to knot theory, traditionally considers all possible embeddings of an abstract graph, the scenario involving specific realizations of graphs within three-dimensional space, where surfaces must not intersect, remains underexplored. This is despite the setting being most relevant for applications taking place in the 3-dimensional Euclidean world. For example, building a graph-like structure using self-assembly of DNA molecules by origami folding requires first finding an unknotted A-trail through the desired spatial graph [4], [7]. When the graph is eulerian and cellular embedded in a surface, an unknotted A-trail can be found by following the boundary of the disks [4].
Since we do not require the surface to be standardly embedded (e.g. a torus is allowed to be knotted), it is in principle possible to obtain all cellular embeddings of spatial graphs by starting from cellular embeddings of free bouquets of circles, on which subdivisions of edges and splittings of vertices are performed on the surface. However, our goal is to construct a surface for a given spatial graph. Note that rotation systems, although equivalent to cellular embeddings for abstract graphs [9], do not ensure the described surface to be free of self-intersections.
A natural approach to construct a surface for a given spatial graph is to glue together non-intersecting properly embedded disks that are bounded by some cycles of the graph. A sufficient albeit unnecessarily strong condition for the existence of such disks is $k$-flatness of the spatial graph [8]. Indeed, it suffices to find a collection of cycles that cover all edges of the graph such that each edge is covered at most twice, and that bound disks satisfying the property above. The union of the disks and the spatial graph produces a cellular embedding of the spatial graph in the resulting surface. In case that every edge belongs to exactly two different cycles, the embedding is a circular (or strong) embedding. However, finding a collection of such disks is in general an intractable problem, as it includes detecting the unknot, non-trivial links with any number of components, Borromean links, and more complicated entanglements such as nonfree knot- and linkless spatial graphs such as the Kinoshita theta-curve.
We therefore follow another approach to construct the surface. Without loss of generality, we can restrict to connected spatial graphs. Indeed, if distinct connected components are nontrivially linked, the spatial graph cannot be cellular embedded, and unlinked components can be considered independently. The idea is to start with a subgraph of the spatial graph that cellular embeds in the sphere, and then iteratively attach handles on which the remaining parts of the graph sit on. If the graph follows a longitude and a meridian of every handle, the resulting embedding is cellular. Since in our construction a longitude of the cylinder that generates a new handle is always followed by some subgraph, cellularity is implied if the two boundary components of the cylinder are attached to different disks of the already generated surface. If the two boundary components of the cylinder are attached to the same disk, it must be ensured that the graph also runs through the meridian of the resulting handle. Furthermore, conditions need to be found that guarantee that the construction results in an embedded surface, i.e., handles should be attached in such a way that they do not intersect the already constructed surfasce parts. For such a procedure to work, we require the spatial graph to have some properties and we call such spatial graphs leveled. They can be viewed as a generalization of planar graphs and are distinct from but related to open book embeddings [2].
We show in proposition 1 that all leveled spatial graphs with up to four levels can be cellular embedded in a surface that itself is embedded in $\mathbb{R}^{3}$ by presenting a constructive argument. We give a combinatorial description for hamiltonian leveled spatial graphs which allows to describe an algorithm whose successful termination constructs a cellular embedding of the spatial graph in algorithm 1. We extend the algorithm to non-hamiltonian leveled graphs in algorithm 2, and finally generalize it to multi-leveled spatial graphs, that are compositions of several leveled spatial graphs.
We give examples where the algorithm fails to find a cellular embedding for a given starting position. However, in all such examples another initiation can be chosen such that the algorithm successfully terminates. We conjecture that this is always the case.
2 Preliminaries
All graphs $G$ are undirected finite graphs. Graphs are allowed to have loops and parallel edges. The fragments (or bridges) of a graph $G$ with respect to a cycle $C$ of $G$ are the closures of the connected components of $G-C$ [10]. Two fragments are conflicting if they either have pairs of endpoints on $C$ that alternate, or have at least three endpoints in common [11]. The conflict graph of a graph $G$ with respect to a cycle $C$ is the graph that has a vertex for each fragment of $G$ with respect to $C$, and two vertices are adjacent if and only if the corresponding fragments conflict. A connected graph is hamiltonian if it contains a cycle that visits every vertex exactly once, i.e., all its fragments with respect to the cycle are the closures of edges or loops. A graph embedding is an embedding $f:G\rightarrow S^{3}$ of a graph $G$ in $S^{3}$ up to ambient isotopy. The corresponding spatial graph $\mathcal{G}$ is the image of this embedding. A spatial graph is trivial if it embeds in $S^{2}$. An abstract graph is planar if it admits a trivial embedding. A diagram of a spatial graph $\mathcal{G}$ is the image of a projection of $\mathcal{G}$ onto $\mathbb{R}^{2}$ such that all intersections in the image are transversal, no vertices and at most two points are mapped to a crossing, and each double point is assigned as an over- or under-crossing. We focus on a special class of spatial graphs, the leveled graphs:
Definition 1.
A leveled spatial graph $\mathcal{G}$ is a connected spatial graph that contains an unknotted cycle $C$ called its spine such that each fragment $f$ of $\mathcal{G}$ with respect to $C$ can be embedded in a disk $D_{f}$ whose boundary is identified with $C$, with interior $\mathring{D_{f}}$ disjoint from $\mathcal{G}-f$ and disjoint from any other $D_{g}$, for $g$ another fragment of $\mathcal{G}$.
We define an ordering of the fragments of $\mathcal{G}$ as follows: Fix a diagram $D$ of a leveled spatial graph $\mathcal{G}$ such that the spine $C$ of $\mathcal{G}$ does not have crossings on $D$. If the fragment $f$ crosses over fragment $g$, then $f>g$. Fragments that are not comparable by the partial order are said to be on the same level. Note that conflicting fragments must be on different levels.
This definition allows to assign the fragments of a leveled spatial graph to their level. The fragments that do not cross over any other fragments are at level $1$, denoted $L_{1}$.
Remark 1.
If the leveled spatial graph $\mathcal{G}$ has $n$ levels $L_{i},1\leq i\leq n$, there is an action of the cyclic group $C_{n}=\langle x\mid x^{n}=1\rangle$ on the partition of levels =$\mathcal{L}=\{L_{1},\dots,L_{n}\}$. The action is given by sending $x\cdot L_{i}$ to $L_{i+1(\mathrm{mod}\ n)}.$
We say that a leveled spatial graph $\mathcal{G}$ is in representative leveling if for each fragment $f$ at level $i>1$, there exists a fragment $g$ at level $i-1$ such that $f$ and $g$ conflict. In other words, $\mathcal{G}$ is in representative leveling if every fragment belongs to the lowest possible level. Every leveled spatial graph can be put into representative leveling by assigning fragments to the lowest possible level. Note that the cyclic reordering of levels does not maintain representative leveling.
Every connected trivial spatial graph is a leveled graph with exactly one level. Consequently, all planar graphs admit a leveled embedding, but this is not true for all graphs. For example, the graph $K_{5}\_K_{5}$ that is obtained by joining two copies of $K_{5}$ by an edge (fig. 11) cannot be embedded as a leveled graph, since every choice of a cycle implies a nonplanar fragment. Therefore, no spine can be found for $K_{5}\_K_{5}$.
In this paper, all surfaces are considered to be closed, orientable, and embedded in $S^{3}$. The spatial graph $\mathcal{G}$ embeds cellular in the surface $\mathcal{S}$ if the complement $\mathcal{S}\setminus\mathcal{G}$ of the spatial graph in the surface is a union of open disks. Two pairs of spatial graphs on surfaces $(\mathcal{G}_{1},\mathcal{S}_{1}),(\mathcal{G}_{2},\mathcal{S}_{2})$ are equivalent if there is an ambient isotopy of $S^{3}$ between $(\mathcal{G}_{1},\mathcal{S}_{1})$ and $(\mathcal{G}_{2},\mathcal{S}_{2})$ that keeps the graphs on the surfaces throughout the ambient isotopy. Note that this notion of equivalence for graph embeddings in surfaces is stronger than the one used in topological graph theory, which only requires surface homeomorphism instead of ambient isotopy.
Without loss of generality, all spatial graphs considered in this paper are connected, since a disconnected graph $\mathcal{G}$ is cellular embedded in a surface $\mathcal{S}$ if and only if each component of $\mathcal{G}$ is cellular embedded in a distinct component of $\mathcal{S}$ and both $\mathcal{G}$ and $\mathcal{S}$ have the same number of components. Indeed, if two connected components $\mathcal{G}^{\prime}$ and $\tilde{\mathcal{G}}$ of $\mathcal{G}$ are embedded in a connected surface $\mathcal{S}$, there exists a simple closed curve $c\in\mathcal{S}\setminus\mathcal{G}$ which encloses $\mathcal{G}^{\prime}$. Since $c$ does not bound a disk in $\mathcal{S}\setminus\mathcal{G}$, the embedding is not cellular.
Since the subgraph of a leveled graph that consists of its spine and fragments of two levels $i$ and $j$ with $i<j$ is trivial by definition, it is always possible to embed a sphere $\mathcal{S}_{0}$ in $\mathbb{R}^{3}$ such that the spine $C$ of $\mathcal{G}$ is embedded in $\mathcal{S}_{0}$, all fragments of level $i$ are embedded in one of the two closed disks $\overline{\mathcal{S}_{0}\setminus C}$, and the fragments of level $j$ on the other disk. The disk on which fragments of level $i$ are embedded in is called the lower hemisphere $\mathcal{S}_{l}$ of $\mathcal{S}_{0}$, and the other disk is called the upper hemisphere $\mathcal{S}_{u}$.
Lemma 1.
Let $\mathcal{G}$ be a spatial graph with a vertex $v$ that is the endpoint of two fragments $f_{1},f_{2}$ with respect to a cycle $C$. It is always possible to split a vertex $v$ into two vertices $v_{1}$ and $v_{2}$ by adding an edge to $C$ such that $v_{1}$ is an endpoint of $f_{1}$ and $v_{2}$ is an endpoint of $f_{2}$, and the fragments in the new graph conflict if and only if they conflict in $\mathcal{G}$. Furthermore, the splitting can be chosen such that for any diagram of $\mathcal{G}$ the crossings remain unchanged.
Proof.
A vertex $v$ of degree $k$ can always be split by replacing a small disk around $v$ with a new disk that leaves the intersection of the graph with the boundary of the disk unchanged, and replaces the vertex $v$ with two vertices $v_{1}$ and $v_{2}$ of degree $m$ and $n$ respectively that are joined by exactly one edge such that $m+n=k+2$. This procedure leaves the crossings of the diagram unchanged. In the following, every vertex splitting is performed in this manner.
Two fragments that share an endpoint are conflicting by definition if they either share two more endpoints, or have three more endpoints that are alternating between the fragments. In the first case, splitting $v$ results in a sequence of alternating endpoints and therefore the fragments remain conflicting. In the second case, splitting $v$ results in four vertices that alternate between being endpoints of $f_{1}$ and $f_{2}$, therefore keeping $f_{1}$ and $f_{2}$ conflicting.
If the two fragments that have the vertex $v$ in common are non-conflicting, following $C$ starting from $v$ runs through all endpoints of one fragment before running through all endpoints of the other. Splitting $v$ as described at above does not introduce an alternation of the endpoints, and therefore keeps the fragments non-conflicting.
By repeatedly splitting vertices on $C$, their degree can be reduced to three. ∎
A hamiltonian leveled spatial graph $\mathcal{H}$ is a leveled spatial graph whose spine is a hamiltonian cycle of the underlying abstract graph. For constructing a cellular embedding for a hamiltonian leveled spatial graph, due to lemma 1, it is possible to restrict without loss of generality to simple graphs, i.e. graphs without loops and parallel edges. In this case, all fragments are edges with two distinct endpoints. Hamiltonian leveled spatial graphs can be completely described combinatorially by their spine list that encodes the order in which fragments are attached to the spine together with their level. If several fragments share an endpoint, the order in the spine list is chosen such that the alternation of fragment labels is avoided. The spine list is unique up to permutation of the labels of parallel edges and cyclic permutations of its symbols, which corresponds to different choices of start points on the spine, see fig. 1.
Figure 1: A hamiltonian leveled spatial graph with spine list $(1^{1},2^{2},3^{1},3^{2},2^{1},3^{2},3^{1},1^{2},2^{1},2^{2},2^{3},1^{2},2^{3},1^{1})$. The superscripts distinguish fragments at same levels.
For general leveled spatial graphs, spine lists descriptions can be produced analogously, introducing extra choices if the choices of splitting common vertices lead to the same minimal number of conflicts between the fragments, see lemma 1. The spine list then determines the leveled graph up to replacing fragments with other fragments that have same endpoints.
Note that the spine list of a leveled embedding $\mathcal{G}$ contains the information of whether or not fragments are conflicting.
Lemma 1 explains how to construct the spine list of a leveled embedding that has fragments with shared endpoints. When a graph is embedded in a surface, contracting edges and splitting vertices does not change the surface. Therefore we can from now on assume without loss of generality that leveled spatial graphs have vertices of degree three on their spine, which allows the use of spine lists.
3 Cellular embeddings of leveled spatial graphs with few levels
In case a leveled spatial graph has a low number of levels, we give an exhaustive solution for its cellular embeddability.
Proposition 1.
If a spatial graph $\mathcal{G}$ has a leveled embedding with at most four levels, then there exists an oriented closed surface $\mathcal{S}$ embedded in $\mathbb{R}^{3}$ where $\mathcal{G}$ cellular embeds.
Proof.
Assume $\mathcal{G}$ is in representative leveling. Let $L_{i}$ be the union of all fragments of level $i$, and let $L_{1,2}:=C\cup L_{1}\cup L_{2}$ be the subgraph of $\mathcal{G}$ that consists of the spine $C$ of $\mathcal{G}$, the fragments of level $1$ and the fragments of level $2$. As $L_{1,2}$ is trivial by definition, there exists an embedded sphere $\mathcal{S}_{0}$ where $L_{1,2}$ is embedded. Since embeddings of connected spatial graphs in a sphere are cellular, the resulting embedding $(L_{1,2},\mathcal{S}_{0})$ is cellular.
Now handles are attached to the upper hemisphere $\mathcal{S}_{u}$ of $\mathcal{S}_{0}$ to construct a surface on which $L_{1,2}\cup L_{3}$ is embedded in. For that, consider the set of disks $D_{u}$ that are the connected components of $\mathcal{S}_{u}\setminus\mathcal{G}$. Since each endpoint $p$ of a fragment in $L_{3}$ lies on the boundary $\partial D_{p}$ of at least one disk $D_{p}\in D_{u}$, an assignment $a(p)=D_{p}\in D_{u}$ with $p\in\partial D_{p}$ can be chosen.
The assignment $a$ induces a clustering of $D_{u}$ in which two disks $D_{p},D_{p^{\prime}}$ belong to the same cluster $Cl_{i}$ if and only if there exists a fragment in $L_{3}$ with two endpoints $p$ and $p^{\prime}$ such that $a(p)=D_{p}$ and $a(p^{\prime})=D_{p^{\prime}}$.
Denote by $F_{i}$ the set of fragments of $L_{3}$ with endpoints in $Cl_{i}$. Since $F_{i}$ is trivial, there is a sphere $S_{F_{i}}$ on which $F_{i}$ embeds in. For each set of endpoints $p_{k}$ of $F_{i}$ for which $a(p_{k})=D_{p_{k}}$, remove a disk $P_{p_{k}}$ from $S_{F_{i}}$ such that $p_{k}\in\partial P_{p_{k}}$ for all $k$ and $q\notin P_{p_{k}}$ for any endpoint $q\neq p_{k}$. This construction yields an $n_{i}$-punctured sphere $\mathcal{S}_{i}$ on which all fragments of $Cl_{i}$ are embedded in.
The embedding of $F_{i}$ in $\mathcal{S}_{i}$ is cellular, because the spatial graph $F_{i}$ together with the boundary cycles of the disks $P_{p_{k}}$ is a connected trivial spatial graph.
Since the punctured spheres $\mathcal{S}_{i}$ are constructed to follow the non-conflicting fragments of level $3$, it is implied that the $\mathcal{S}_{i}$ are not intersecting. Removing the interiors of the disks that are in the image of $a$ from $\mathcal{S}_{u}$ allows to construct an oriented closed non-intersecting surface. This is done by identifying their boundaries with the boundaries of the punctures of $\mathcal{S}_{i}$ that respects the induced orientations on the boundaries of the surfaces.
Analogously, $L_{4}$ is (cellular) embedded in $n_{j}$-punctured punctured spheres $\mathcal{S}_{j}$ that are attached to the lower hemisphere $\mathcal{S}_{l}$.
The punctured spheres on which $L_{4}$ embeds in do not intersect the surface where $L_{1,2}\cup L_{3}$ is embedded. This is due to there being four levels only: each level is attached to only one of the hemispheres and the crossing types between the levels are of only one type, making the attachment of handles to the upper hemisphere and to the lower hemisphere independent of one another. The result is an embedding of $\mathcal{G}=L_{1}\cup L_{2}\cup L_{3}\cup L_{4}$ in a handlebody of genus $\sum_{i}(n_{i}-1)+\sum_{j}(n_{j}-1)$.
It remains to be seen that the embedding is cellular. The embedding of $\mathcal{G}$ in the resulting handlebody is cellular since each $\mathcal{S}_{k}$ is attached to different disks of $\mathcal{S}_{0}\setminus\mathcal{G}$, and $F_{k}$ is cellular embedded in $\mathcal{S}_{k}$, and $L_{1,2}$ is cellular embedded in $\mathcal{S}_{0}$.
Corollary 1.
The genus of the surface $\mathcal{S}$ obtained in proposition 1 depends on the leveled embedding itself and the definition of the function $a(p)$ in the proof. There is a unique $a$ if fragments have endpoints of degree three only. For degrees greater than three, $a$ is uniquely defined if and only if neither fragments of level 3 and fragments of level 2 nor fragments of level 4 and fragments of level 1 have endpoints in common.
The restriction to four levels in proposition 1 assures both the cellularity of the resulting embedding and the absence of self-intersections of the resulting surface. The construction of the proof of proposition 1 ensures neither property for a leveled embedding with five or more levels, as can be seen in the following example.
Example 1.
Consider the leveled embedding $\mathcal{G}$ shown left in fig. 13, that is given by the spine list
$(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},3^{1},3^{2},3^{3},4^{1},4^{2},4^{3},5^{1},5^{2},5^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},3^{3},3^{2},3^{1},4^{3},4^{2},4^{1},5^{3},5^{2},5^{1}).$
We can apply the algorithm in the proof of proposition 1 to embed any four levels in a surface $\mathcal{S}$; without loss of generality choose the levels 1, 2, 4, 5. This leaves level 3 to be embedded. The endpoints of fragments of level 3 lie on the boundary of two disks of $\mathcal{S}\setminus\{L_{1}\cup L_{2}\cup L_{4}\cup L_{5}\}$. Denote by $D_{u}$ (respectively $D_{l}$) the disk which has interior points in common with the upper (resp. lower) hemisphere $S_{u}$ (resp. $S_{l}$) of the sphere that is produced in the first step of constructing $\mathcal{S}$. Extending $\mathcal{S}$ to accommodate $L_{3}$ cellularly requires a single cylinder $P$ to be placed under fragments $3^{1},3^{2},3^{3}$ such that $P$ is attached to $\mathcal{S}$ with one end glued to $D_{u}$ and the other to $D_{l}$. However, the resulting surface is non-orientable. This can be seen as follows: the order in which the three fragments are attached to the spine defines an orientation of the spine. As the three fragments do not conflict in $\mathcal{G}$ with respect to the spine, and $P$ is glued to $D_{u}$ on one end and to $D_{l}$ on the other, the two ends of $P$ are oriented with the same orientation as the spine. This is, if the orientation on one boundary component of the cylinder is clockwise, it is counterclockwise for the other. The boundary orientation does therefore not correspond to the orientation that is obtained when considering the two ends of the cylinder as naturally identified. This implies that the surface obtained by gluing $P$ to the oriented surface $\mathcal{S}$ is non-orientable.
This shows that the surface with $\mathcal{G}$ embedded is either cellular or free of self-intersections, but cannot be both.
4 Constructing the surface
4.1 Placing cylinders under fragments
In this section, given a spatial graph $\mathcal{G}$ embedded in a surface $\mathcal{S}$ formed by gluing cylinders consecutively to a sphere, we derive restrictions on the placement of fragments on the cylinders forming $\mathcal{S}$ such that the embedding of $\mathcal{G}$ in $\mathcal{S}$ is leveled.
Unless specified otherwise, we suppose that the spine of a leveled embedding is a hamiltonian cycle of the underlying graph. Under this assumption, every fragment is an edge of the graph. To denote the spine list of leveled spatial graphs, we use letters as variables for fragments.
Definition 2.
Let $\mathcal{G}$ be a leveled embedding in representative leveling with $n$ levels. We say that $\mathcal{G}$ is consecutively embedded if it is embedded in a surface $\mathcal{S}\cong S^{2}\cup_{\varphi_{1}}P_{1}\cup_{\varphi_{2}}\dots\cup_{\varphi_{m}}P_{m}$ that is built successively by gluing the ends of a cylinder $P_{i}$ to $\mathcal{S}_{i-1}=S^{2}\cup_{\varphi_{1}}P_{1}\cup_{\varphi_{2}}\dots\cup_{\varphi_{i-1}}P_{i-1}$ by a map $\varphi_{i}$ such that:
  • the spine of $\mathcal{G}$ is embedded as equator of $\mathcal{S}_{0}=S^{2}$ and levels 1 and 2 are embedded respectively in the lower and upper hemisphere of $S^{2}$;
  • each cylinder $P_{i}$ carries at least one fragment of $\mathcal{G}$;
  • the subgraph of $\mathcal{G}$ embedded in $\mathcal{S}_{i}$ is called $\mathcal{G}_{i}$;
  • $\varphi_{i}$ glues $P_{i}$ to $\mathcal{S}_{i-1}$ such that the two ends of $P_{i}$ are glued to two subsets of one or two disks of $\mathcal{S}_{i-1}\setminus\mathcal{G}_{i-1}$ which are each homeomorphic to a disk and incident to the spine of $\mathcal{G}$.
Definition 3.
Let $\mathcal{G}$ be a leveled embedding in representative leveling with $n$ levels, consecutively embedded in a surface $\mathcal{S}_{i}$. Let $P_{i}$ be a cylinder which is glued to two not necessarily distinct disks $D_{1},D_{2}$, such that $\mathcal{S}_{i}=\mathcal{S}_{i-1}\cup_{\varphi_{i}}P_{i}$. A subset of the spine list of $\mathcal{G}$ consisting of the endpoints of the fragments embedded in $P_{i}$ which lie on one end of $P_{i}$ is called a spine sublist of $P_{i}$. Reversing the order of the endpoints in one of the spine sublist results in the disk sublists of $P_{i}$. This is, to obtain the disk sublists, the boundary components of the cylinder are considered as naturally identified.
Remark 2.
The embedding of fragments $f_{1},\dots,f_{n}$ of a leveled spatial graph $\mathcal{G}$ in a cylinder $P$ where every fragment has endpoints lying on the boundary of $P$ is cellular in $P$ if and only if at least one fragment $f_{i}$ has endpoints on both ends of $P$.
Indeed, $f_{i}$ cuts $P$ along the longitude, therefore creating a disk. Any other fragment $f_{j}$ then cuts a disk into two disks. Therefore, the resulting embedding $(f_{1},\dots,f_{n},P)$ is cellular. On the other hand, if the embedding $(f_{1},\dots,f_{n},P)$ is cellular, every simple closed curve of $P$ parallel to the boundary of $P$ (a meridian of $P$) must intersect at least one fragment $f_{i}$. Therefore, there exists a fragment $f_{i}$ with endpoints on both ends of $P$.
Lemma 2.
Let $\mathcal{G}$ be a leveled spatial graph consecutively embedded in a surface $\mathcal{S}$. Let $P_{i}$ be a cylinder which is glued to two (not necessarily distinct) disks $D_{1},D_{2}$ of $\mathcal{S}_{i-1}\setminus\mathcal{G}_{i-1}$. If a fragment $f$ embedded on $P_{i}$ has two or more halftwists with respect to $P_{i}$, then no other fragment with endpoints on both ends of $P_{i}$ has more than one halftwist with respect to $P_{i}$. Moreover, if $f$ has three or more halftwists with respect to $P_{i}$, then $f$ is the only fragment with endpoints on both ends of $P_{i}$.
Proof.
If there exists a pair of fragments with endpoints on both ends of $P_{i}$ with two or more halftwists with respect to $P_{i}$, then the two fragments need to cross over and under each other. Therefore, the embedding would not be leveled. Moreover, if $f$ has three or more halftwists with respect to $P_{i}$, then any other fragment with endpoints on both ends of $P_{i}$ needs to be parallel to it on $P_{i}$. Therefore, the difference of halftwists with respect to $P_{i}$ of the two fragments is at most 1. This is impossible due to the previous argument. ∎
Proposition 2.
Let $\mathcal{G}$ be a leveled spatial graph consecutively embedded in a surface $\mathcal{S}$. Let $P_{i}$ be a cylinder which is glued to two (not necessarily distinct) disks $D_{1},D_{2}$ of $\mathcal{S}_{i-1}\setminus\mathcal{G}_{i-1}$. Then the disk sublists of $P_{i}$ must be one of the following:
  1. 1.
    $(a_{1},\dots,a_{m},b_{1},\dots,b_{p},c,a_{m},\dots,a_{1}/d_{1},\dots,d_{q},c,b_{1},\dots,b_{p},d_{q},\dots,d_{1});$
  2. 2.
    $(a_{1},\dots,a_{m},c,b_{1},\dots,b_{p},a_{m},\dots,a_{1}/d_{1},\dots,d_{q},b_{1},\dots,b_{p},c,d_{q},\dots,d_{1});$
  3. 3.
    $(s_{1},\dots,s_{n},\alpha_{1},\dots\alpha_{k},\beta_{1},\dots,\beta_{h},s_{n},\dots,s_{1}/t_{1},\dots,t_{q},\beta_{1},\dots,\beta_{h},\alpha_{1},\dots,\alpha_{k},t_{q},\dots,t_{1}),$
  4. 4.
    $(a_{1},\dots,a_{m},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p},\beta_{1},\dots,\beta_{h},a_{m},\dots,a_{1}/\beta_{1},\dots,\beta_{h},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p});$
  5. 5.
    $(a_{1},\dots,a_{m},c,\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p},a_{m},\dots,a_{1}/\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p},c);$
  6. 6.
    $(a_{1},\dots,a_{m},b_{1},\dots,b_{p},\beta_{1},\dots,\beta_{h},c,a_{m},\dots,a_{1}/\beta_{1},\dots,\beta_{h},c,b_{1},\dots,b_{p});$
where the slash “/” separates the two disk sublists of $P_{i}$, the letters used are variables, $c$ denotes a fragment with at least two halftwists, Greek letters denote fragments with exactly one halftwist, same Latin letters indicate fragments at same level, some letters may be absent except there must be at least one fragment having endpoints on both ends of $P_{i}$, sequences of Latin letters $x_{1},\dots,x_{r},x_{r},\dots,x_{1}$ can be added in any point of the disk sublists.
Proof.
All endpoints of fragments lie on the boundary of $P_{i}$ by construction. The placement of fragments on $P_{i}$ is restricted by how fragments can be placed on a cylinder in general, and secondly by the requirement of leveledness.
For the first part, consider a cylinder $P$ that is not attached to a surface. There are two ways to embed edges in $P$ such that their endpoints lie on the boundary of the cylinder: either the two endpoints of an edge are attached to both ends of the cylinder, or both endpoints are attached to the same end. The endpoints of the fragments embedded in $P_{i}$ can be moved along the connected component of the boundary of $P_{i}$ to which the endpoints belong to while maintaining their embedding in $P_{i}$. Lemma 2 gives conditions for the embedding to be leveled: in order to obtain the lists presented, consider the cylinder $P_{i}$ embedded in $\mathbb{R}^{3}$ but not glued to $\mathcal{S}_{i}$ depicted in fig. 2. Shifting the endpoints of the fragments along the boundary components of $P_{i}$ while respecting the conditions given by lemma 2 yields the six presented lists. ∎
Figure 2: The endpoints of the fragments can be moved along the connected components of the boundary of the cylinder $P$ which is not attached to the surface.
Remark 3.
Lemma 2 and proposition 2 restrict how fragments are embedded in a cylinder $P_{i}$ and the possible disk sublists of $P_{i}$ in a consecutively embedded leveled spatial graph.
Proposition 2 allows to extend proposition 1:
Corollary 2.
If a spatial graph $\mathcal{G}$ admits a leveled embedding with at most five levels such that at least one level consists of at most two edges, then there exists a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds.
Proof.
Let the level which consists of at most two edges be level 3 of the embedding. Perform the algorithm shown in the proof of proposition 1 on the other four levels of $\mathcal{G}$. This creates a surface where four levels of $\mathcal{G}$ are cellular embedded and the embedding has at least two disks which intersect $\mathcal{S}_{0}$ in different hemispheres. If level 3 consists of one fragment or of two fragments whose endpoints lie in the boundaries of one pair of disks, then embed the fragments of level 3 in one cylinder; this is possible due to proposition 2 and the fact that the two fragments do not conflict. Otherwise level 3 consists of two fragment whose endpoints do not lie in the boundaries of one pair of disks. Then each fragment can be placed on a separate cylinder.
Definition 4.
Let $\mathcal{G}$ be a leveled embedding which is consecutively embedded in $\mathcal{S}$. Let $P$ be a cylinder in $\mathcal{S}$. We say that $P$ is an upper cylinder if the fragments embedded in it are at higher level than the fragments embedded in $\mathcal{S}_{0}$. Otherwise, $P$ is called a lower cylinder.
The neighbourhood of the spine of $\mathcal{G}$ in the surface $\mathcal{S}$ has two components. We call them $N_{u}$ and $N_{l}$. For an endpoint $p$ of a fragment $f$ embedded in a cylinder $P$, let $N_{p}$ be a small neighbourhood of $p$ in $P$. Then $N_{p}$ is contained in exactly one of $N_{u}$ or $N_{l}$. We say that $p$ is an upper endpoint (resp. lower endpoint) if $N_{p}$ is contained in $N_{u}$ (resp. $N_{l}$). (fig. 3)
Figure 3: An upper cylinder $P$ with a fragment $f$ with upper endpoints and a lower cylinder $P^{\prime}$ with a fragment $g$ with upper endpoints.
Figure 4: The situation described in cases 1 and 2 of lemma 3.
Lemma 3.
Consider a cylinder $P_{i}$ of the consecutively embedding of $\mathcal{G}$ in $\mathcal{S}.$
  1. 1.
    If $P_{i}$ is an upper cylinder, the fragments of $P_{i}$ with endpoints on only one end of $P_{i}$ and with at least one upper endpoint must be at lower level than the fragments they conflict with that have endpoints on both ends of $P_{i}$ and at least one upper endpoint in the spine sublist that contains the conflict.
  2. 2.
    If $P_{i}$ is an upper cylinder and has a fragment with upper endpoints on both ends of $P_{i}$ and two halftwists with respect to $P_{i}$, then that fragment must be at lower level than the other fragments with endpoints on both ends of $P_{i}$ and at least one upper endpoint.
  3. 3.
    If $P_{i}$ is a lower cylinder, the fragments of $P_{i}$ with endpoints on only one end of $P_{i}$ and with at least one upper endpoint must be at lower level than the fragments they conflict with that have endpoints on both ends of $P_{i}$ and at least one upper endpoint in the spine sublist that contains the conflict.
  4. 4.
    If $P_{i}$ is a lower cylinder and has a fragment with upper endpoints on both ends of $P_{i}$ and two halftwists with respect to $P_{i}$ must be at lower level than the other fragments with endpoints on both ends of $P_{i}$ and at least one upper endpoint.
  5. 5.
    The previous four statements with “lower level” replaced by “higher level” are true if the fragments mentioned have at least one lower endpoint.
Proof.
In the first case, let an upper endpoint of the fragment $g$ with endpoints on both ends of $P_{i}$ lie in between the two upper endpoints of a fragment $f$ with endpoints on only one end of $P_{i}$. Then $f$ must be at lower level than $g$ by construction of $P_{i}$ being an upper cylinder.
In the second case, if $P_{i}$ has a fragment $f^{\prime}$ with two halftwists with respect to $P_{i}$ and an upper endpoint, then any other fragment attached to both ends of $P_{i}$ and having an upper endpoint has at most one halftwist with respect to $P_{i}$ and hence is at higher level than $f^{\prime}$.
Case 3 follows from case 1, and similarly case 4 follows from case 2 by applying a reflection with respect to the plane where the spine lies. Case 5 follows from cases 1 to 4 since if a fragment has a lower endpoint, then it has one more halftwist with respect to $P_{i}$ than a fragment with an upper endpoint.
Proposition 3.
Suppose fragments $a$ and $b$, with $a$ at lower level than $b$ and both with at least one upper endpoint, are embedded in an upper cylinder $P$, with the disk sublists of $P$ restricted to $a$ and $b$ being $(a,b/a,b)$ or $(a,b,a/b)$. Suppose there exists a fragment $x$ such that an upper endpoint of $x$ lies on the spine between a pair of endpoints of $(a,b)$ in one disk sublist of $P$ (i.e. the disk sublists restricted to $a,b$ and such endpoint of $x$ are $(a,x,b/a,b)$, $(a,x,b,a/b)$ or $(a,b,x,a/b)$). If $x$ is at same or lower level than $a$, then $x$ is not embedded in an upper cylinder.
Proof.
We only prove the case where the disk sublists restricted to $a,b$ and the endpoint of $x$ lying on the spine between the pair of endpoints of $(a,b)$ as in the statement is of the form $(a,x,b/a,b)$. The other cases are similar. We prove the counterpositive: if $x$ is embedded in an upper regular cylinder, then $x$ is at higher level than $a$. (fig. 5)
If $x$ is embedded in $P$, then it must be at higher level than $a$, by lemma 3.
Consider the case where $x$ is embedded in another upper cylinder $\tilde{P}$. The upper endpoint of $x$ in between the pair of endpoints of $(a,b)$ lies on the intersection between the spine and the disk to which $P$ is glued to. Also, $P$ and $\tilde{P}$ are upper cylinders. Therefore $\tilde{P}$ crosses over $P$, which implies that $x$ crosses over $a$ or $b$. ∎
Figure 5: Illustration of the counterpositive statement of proposition 3.
Corollary 3.
Let $a$ and $b$ be fragments embedded in a cylinder $P$ with $a$ at lower level than $b$. Let $x$ be another fragment embedded in a cylinder $P^{\prime}$ (not necessarily distinct from $P$) which has an endpoint lying on the spine between a pair of endpoints of $(a,b)$ in one disk sublist of $P$.
  1. 1.
    Assume $P$ is a lower cylinder and $a,b,x$ are lower endpoints. If $x$ is at same or higher level than $b$, then $x$ is not embedded in a lower cylinder.
  2. 2.
    Assume $P$ is an upper cylinder and $a,b,x$ are lower endpoints. If $x$ is at same or higher level than $a$, then $x$ is not embedded in an upper cylinder.
  3. 3.
    Assume $P$ is a lower cylinder and $a,b,x$ are upper endpoints. If $x$ is at same or lower level than $b$, then $x$ is not embedded in a lower cylinder.
Figure 6: Illustration of the counterpositive statement of case 2 of corollary 3. Notice that one end of $P^{\prime}$ is glued to $P$.
Proof.
We only need to prove 2, because 1 follows from proposition 3 and 3 follows from 2 by applying a reflection with respect to the plane where the spine lies. We prove the counterpositive: if $P$ and $P^{\prime}$ are upper cylinders and the endpoints of $a,b,x$ are lower endpoints, then $x$ is at lower level than $a$ (fig. 6). Let $P$ be an upper cylinder and $x$ have a lower endpoint in between two lower endpoints of $a$ and $b$ in the same disk sublist. If $x$ can also be embedded in $P$, then $x$ is at lower level than $b$ due to lemma 3. If $x$ is embedded in another upper consonant cylinder $P^{\prime}$, then it conflicts with either both or exactly one of $a$ and $b$. In the latter case, the crossing number of $x$ with the fragment it does not conflict with must be 0. If $x$ conflicts with $a$, then $x$ is at lower level than $a$, because else $x$ would need to pierce the cylinder $P$ or cross over $P$, making it cross also fragment $b$; contradicting $x$ having crossing number 0 with $b$.
If $x$ conflicts only with $b$, it is not possible that $x$ is at higher level than $b$, because then $P^{\prime}$ would cross over $P$, which implies that $x$ crosses over $a$ non-trivially. This is a contradiction of the fact that $x$ has crossing number 0 with $a$.
The case where $x$ conflicts with both $a$ and $b$ implies that $x$ is at lower level than $a$. Indeed, it is not possible that $x$ is at a level in between the levels of $a$ and $b$, because $P^{\prime}$ would need to pierce through $P$. If $x$ is at higher level than $b$, then $P^{\prime}$ must cross over $P$. But this is not possible because $x$ has a lower endpoint in between two lower endpoints of $a$ and $b$. Therefore, $x$ needs first to cross under $a$ or $b$ before crossing over $b$, which is a contradiction to being leveled. Hence in this case $x$ is at lower level than $a$. ∎
4.2 Attaching cylinders to the surface
4.2.1 Conditions to avoid self-intersections of the surface
In this section, given a hamiltonian leveled spatial graph $\mathcal{G}$, we investigate how to produce a surface $\mathcal{S}$ where $\mathcal{G}$ embeds such that $\mathcal{S}$ has no self-intersections.
We examine when gluing a cylinder $P$ to $\mathcal{S}$ causes self-intersections. For that, we need to consider its interactions with already attached cylinders having endpoints in their spine sublist which, with respect to the spine list of $\mathcal{G}$, alternate with, are contained by, or contain endpoints in a spine sublist of $P$. Up to reflection of $\mathcal{S}$ with respect to the plane where the spine lies, four cases can arise at every endpoint:
  1. 1.
    upper cylinder with an upper endpoint and lower cylinder with a lower endpoint,
  2. 2.
    upper cylinder with an upper endpoint and upper cylinder with a lower endpoint,
  3. 3.
    upper cylinder with an upper endpoint and lower cylinder with an upper endpoint,
  4. 4.
    upper cylinder with a lower endpoint and lower cylinder with an upper endpoint.
We need to consider if the cylinders intersect each other only at the ends where their endpoints lie. If the cylinders do not intersect there, then it is possible to embed them without intersections.
We assume that we have a hamiltonian leveled spatial graph $\mathcal{G}$ consecutively embedded on a surface $\mathcal{S}$.
Definition 5.
A cylinder $P$ and the endpoints of the fragments embedded in it interact with another cylinder $P^{\prime}$ and its endpoints if there exists an endpoint of a fragment of $P^{\prime}$ which, in the spine list of $\mathcal{G}$, is in between two endpoints belonging to the same spine sublist of $P$ or vice versa. In particular, we say that $P$ encloses $P^{\prime}$ if a spine sublist of $P^{\prime}$ is entirely in between two endpoints belonging to the same spine sublist of $P$. If a spine sublist of $P$ interacts with a spine sublist of $P^{\prime}$ but neither encloses the other, then the endpoints of their fragments alternate.
In case 1, the upper cylinder $P$ has an upper endpoint $e$ while the lower cylinder $P^{\prime}$ has a lower endpoint $e^{\prime}$. The ends of the two cylinders where $e$ and $e^{\prime}$ lie do not intersect each other as the plane where the spine lies separates them. (fig. 7)
We cover cases 2, 3 and 4 in proposition 4, proposition 5 and corollary 4 respectively.
Figure 7: An upper cylinder $P$ with upper endpoints can freely interact with a lower cylinder $P^{\prime}$ with lower endpoints.
Proposition 4.
Let $P$ and $P^{\prime}$ be upper cylinders. If the fragments $f_{1},\dots,f_{s}$ of $P$ which conflict with a fragment of $g_{1},\dots,g_{t}$ of $P^{\prime}$ are at lower level than the fragments of $P^{\prime}$ they conflict with, then the lower endpoints of $f_{1},\dots,f_{s}$ do not interact with the upper endpoints of $g_{1},\dots,g_{t}$ of $P^{\prime}$ they conflict with. (See fig. 8 for the counterpositive.)
Proof.
We prove the case where $P^{\prime}$ encloses $P$, arguing by contradiction; the other cases are similar. Suppose that the fragments $f_{1},f_{2}$ of $P$ conflict with fragment $g_{2}$ of $P^{\prime}$ and not with $g_{1}$. Moreover, suppose that the upper endpoints of $g_{1}$ and $g_{2}$ enclose the lower endpoints of $f_{1},f_{2}$.
$f_{1}$ and $f_{2}$ are at lower level than $g_{2}$. Therefore, if $P^{\prime}$ encloses $P$, then $P$ needs to cross under $P^{\prime}$, because $f_{1},f_{2}$ have lower endpoints enclosed by upper endpoints of $g_{1},g_{2}$. This implies $f_{1},f_{2}$ cross under $g_{1}$, which contradicts $f_{1}$ and $f_{2}$ not conflicting with $g_{1}$. ∎
Figure 8: Illustration of the counterpositive statement of proposition 4.
Proposition 5.
Let $P$ be a lower cylinder and $P^{\prime}$ be an upper cylinder such that $P=P_{j}$ and $P^{\prime}=P_{i}$ following the notation of definition 2. If $i>j$, then there does not exist a fragment $f$ of $P^{\prime}$ which has an upper endpoint lying on the spine list in between two upper endpoints of fragments $g_{1},g_{2}$ of $P$ belonging to the same spine sublist of $P$.If $i<j$, then $P$ and $P^{\prime}$ do not intersect each other if and only if there exists a path from $g_{1}$ to $g_{2}$ along the boundary of the disk where $g_{1},g_{2}$ lie which only crosses over $P^{\prime}$ or below the fragments embedded in $\mathcal{S}_{0}$.
Proof.
We prove the counterpositive statement for the case $i>j$.
By hypothesis the spine list restricted to the endpoints of the fragments $f,g_{1},g_{2}$ on the interacting ends of $P$ and $P^{\prime}$ is of the form $(g_{1},f,g_{2})$. If $i>j$, then the endpoint of $f$ in between those of $g_{1}$ and $g_{2}$ is contained in the boundary of $P$. Being $P$ a lower cylinder, this means that $f$ needs to cross under $g_{1}$ or $g_{2}$, which contradicts $P^{\prime}$ being an upper cylinder.
In case $i<j$, if there exists a path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ along the boundary of the disk with endpoints of $g_{1},g_{2}$ on its boundary which only crosses over $P^{\prime}$ or below the fragments embedded in $\mathcal{S}_{0}$, then it is possible to slightly move the path in order to form a disk $D$ to which the cylinder $P^{\prime}$ can be glued. As the disk only crosses over $P^{\prime}$ or below the fragments embedded in $\mathcal{S}_{0}$, then $P$ and $P^{\prime}$ do not intersect.
If instead there is no path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ along the boundary of the disk with endpoints of $g_{1},g_{2}$ on its boundary which only crosses over $P^{\prime}$ or below the fragments embedded in $\mathcal{S}_{0}$, meaning that every path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ crosses under $P^{\prime}$ and above the fragments embedded in $\mathcal{S}_{0}$, then $P$ and $P^{\prime}$ intersect by the intermediate value theorem applied to any path and fragment $f$. ∎
Corollary 4.
Let $P$ be a lower cylinder and $P^{\prime}$ be an upper one such that $P$ has an upper endpoint $e$ and $P^{\prime}$ has a lower endpoint $e^{\prime}$. Then the two cylinders can be placed on the resulting surface as long as they are not interacting.
Proof.
Follows from proposition 5. By symmetry, neither cylinder can enclose the other nor alternate with the other. ∎
4.2.2 Conditions to ensure cellularity
In this section, we investigate which hamiltonian leveled spatial graphs that are consecutively embedded are cellular.
Lemma 4.
Let $\mathcal{G}_{i}$ be a hamiltonian leveled spatial graph consecutively embedded in a surface $\mathcal{S}_{i}.$ Suppose fragments $f_{1},\dots,f_{t}$ of the spatial graph $\mathcal{G}_{i}$ are embedded in a cylinder $P_{i}$, while the spatial subgraph $\mathcal{G}_{i-1}\subset\mathcal{G}_{i}$ is cellular embedded in the surface $\mathcal{S}_{i-1}$ and $\mathcal{G}_{i}-\mathcal{G}_{i-1}=\{f_{1},\dots,f_{t}\}.$
  1. (1)
    If $P_{i}$ is glued to two different disks $D_{1},D_{2}$ of the cellular embedding of $\mathcal{S}_{i-1}\setminus\mathcal{G}_{i-1}$, then the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular.
  2. (2)
    If $P_{i}$ is glued to a single disk $D$, then the resulting embedding is cellular if and only if there exists at least one conflicting pair of fragments in $f_{1},\dots f_{t}$ in the conflict graph with respect to the boundary of $D.$
Proof.
Assume that the fragment which is attached to both ends of $P_{i}$ is $f_{j}$.
  1. 1.
    In the case of $P_{i}$ being glued to two different disks, consider a closed path $\gamma_{1}\in\mathcal{G}$ formed by $f_{j}$ and part of the spine joining the endpoints of $f_{j}$. All simple closed curves on $\mathcal{S}_{i}\setminus\mathcal{G}_{i-1}$ which do not bound a disk in $\mathcal{S}_{i}\setminus\mathcal{G}_{i-1}$ intersect $\gamma_{1}$ or the boundaries of $D_{1}$ or $D_{2}.$ This proves that gluing $P_{i}$ keeps the embedding cellular.
  2. 2.
    Consider $P_{i}$ glued to $D$ as a torus $T$ with a marked point $t_{0}$, which coincides with the boundary of $D$. Two loops embedded in a torus form a cellular embedding if and only if their crossing number is greater than 0. In our case, loops are fragments and two fragments can cross each other only in $t_{0}$. Therefore, $((T,t_{o}),P_{i})$ is cellular if and only if there exist at least two fragments $f_{1},f_{2}$ such that their endpoints alternate on the boundary of $D.$
Example 2.
Consider the leveled spatial graph $\mathcal{G}$ given by $(3,1,4,2,1,3,4,2).$ We would like to embed $\mathcal{G}$ consecutively such that all endpoints are upper endpoints, fragment $2$ is embedded in the upper hemisphere of a sphere $\mathcal{S}_{0}$, fragment $1$ is embedded in a lower cylinder $P_{1}$ glued to the two disks formed by fragment $2$ (notice that this merges the two disks of the upper hemisphere into a single disk $D$) and fragments $3$ and $4$ are embedded together in an upper cylinder $P_{2}$ glued to $D$. See fig. 9 for a picture of the embedding.
This embedding of $\mathcal{G}$ in $\mathcal{S}_{2}$ is not cellular, because fragments $3$ and $4$ do not conflict with respect to the boundary of $D$ by lemma 4. Note in particular that $3$ and $4$ conflict with respect to the spine of $\mathcal{G}$, but this does not make the embedding cellular.
Figure 9: The embedding described in example 2.
In proposition 6 it is shown that the amount of disks obtained from gluing the cylinder $P_{i}$ to a surface $\mathcal{S}_{i-1}$ depends only on how many fragments are embedded in $P_{i}$, as long as the embedding is cellular.
Proposition 6.
Let $\mathcal{G}_{i}$ be a hamiltonian leveled spatial graph consecutively embedded in a surface $\mathcal{S}_{i}$. Let $\mathcal{G}_{i-1}$ be the (leveled) subgraph of $\mathcal{G}_{i}$ consecutively embedded in $\mathcal{S}_{i-1}$ and $\mathcal{G}_{P_{i}}$ be the subgraph of $\mathcal{G}_{i}$ embedded in $P_{i}$. Suppose that $\mathcal{G}_{i-1}$ is cellular embedded in $\mathcal{S}_{i-1}$ with $n$ disks. Suppose that $\mathcal{G}_{P_{i}}$ is cellular embedded in $P_{i}$ and $t$ is the number of fragments of $\mathcal{G}_{P_{i}}\subset\mathcal{G}_{i}$.
Then, $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $n-2+t$ disks.
Proof.
The cellular embedding of $\mathcal{G}_{P_{i}}$ in $P_{i}$ has $t$ disks. Indeed, the first fragment of $\mathcal{G}_{P_{i}}$ which is attached to both ends of $P_{i}$ cuts an annulus into a disk, and every other fragment of $\mathcal{G}_{P_{i}}$ cuts an existing disk in two disks, adding one disk to the total per fragment.
If $P_{i}$ is glued to two different disks of $\mathcal{S}_{i-1}$, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $n-2+t$ disks.
If $P_{i}$ is glued to a single disk $D$ and the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular as seen in lemma 4, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $(n-1)+(t-1)$ disks. Indeed, for each end of $P_{i}$, the disk of $P_{i}\setminus\mathcal{G}_{P_{i}}$ that contains $\partial P_{i}$ which is not part of the spine is merged with $D$. Notice that, due to cellularity, the disks merged on the two ends of $P_{i}$ are different. ∎
Given that the embedding is cellular, we adapt the “Face tracing algorithm” of a cellular graph embedding explained in [5] to be able to trace the boundaries of its disks.
Fact 1 (Face tracing algorithm).
Let $\mathcal{G}$ be a leveled spatial graph with hamiltonian spine cellular consecutively embedded in a surface $\mathcal{S}$. Consider the spine list of $\mathcal{G}$ and an endpoint $e$ of a fragment $f$. Go to the endpoint after $e$ on the spine list of $\mathcal{G}$ with same type as $e$ (i.e. upper if $e$ is an upper endpoint and lower otherwise). Go to the other endpoint of the fragment with endpoint $e$ and repeat this process. The obtained closed walk is the boundary of a disk of the embedding. Starting at different endpoints gives all disks of the cellular embedding.
4.3 The algorithm for the surface construction
In this section, we provide an algorithm which, given as input a hamiltonian leveled spatial graph $\mathcal{G}$ in representative leveling by means of its spine list $\ell$, gives as output a surface $\mathcal{S}$ where $\mathcal{G}$ is embedded cellular and consecutively. The algorithm chooses two consecutive levels that are placed on a sphere and scans through all possible partitions of the remaining set of fragments of $\mathcal{G}$ into subsets, checking whether the fragments in each subset of a partition can be placed on a cylinder $P$ together. Partitions which do not yield a surface are identified using the results of the previous sections and are discarded. The algorithm ends successfully if all subsets of one partition are placed on cylinders and ends unsuccessfully if all partitions are discarded.
The algorithm is extended to leveled spatial graphs with any spine, not necessarily hamiltonian, and to multi-leveled embeddings.
Algorithm 1.
Let $\mathcal{G}$ be a hamiltonian leveled spatial graph with spine $C$, spine list $\ell$ and $n$ levels.
For $1\leq c<n$ choose the levels $c$ and $c+1$. Generate an embedded sphere $\mathcal{S}_{0}=S^{2}$ with the equator being the spine $C$ and with fragments at level $c$ being embedded in the lower hemisphere and fragments at level $c+1$ in the upper hemisphere. Call this spatial subgraph $\mathcal{G}_{0}$.
Let $\mathcal{F}_{-\{c,c+1\}}$ be the set of $N$ fragments of $\mathcal{G}$ that are not in level $c$ and $c+1$. Generate all possible partitions $P_{k},1\leq k\leq\text{e}^{-1}\sum_{m=0}^{\infty}\frac{m^{N}}{m!}$ (Dobińsky’s formula [3]), of $\mathcal{F}_{-\{c,c+1\}}$ into disjoint nonempty subsets. That is, the elements of $P_{k}$ are subsets $P_{k}^{j}$, $1\leq j\leq J(k)$ of fragments, where $J(k)$ is the number of the subsets in $P_{k}$. The fragments in $P_{k}^{j}$ will be candidates for fragments embedded in one cylinder which by abuse of notation we also denote by $P_{k}^{j}$.
For $P_{k}$, there are $2^{J(k)}$ choices of labelling its elements with labels $\{U,L\}$. This label indicates whether the cylinder $P_{k}^{j}$ is an upper or lower cylinder. (See example 3 for the notation.)
The endpoints of the fragments in $P_{k}^{j}$ can be separated into two disk sublists. That is, for $P_{k}^{j}$ with $n$ endpoints, there are $\frac{1}{2}(n-1)n$ separations. Labelling the endpoints with $\{u,l\}$ results in $2^{n-1}(n-1)n$ labelled separations of $P_{k}^{j}$. This label indicates whether an endpoint is an upper or lower endpoint.
Enumerate the partition with labelled cylinders and labelled endpoints as $\mathcal{P}_{i}$. A labelled element in $\mathcal{P}_{i}$ is denoted by $P_{i}^{j}$. (See example 4.)
For $i$ from $1$ to the number of labelled partitions, consider $\mathcal{P}_{i}$. For $0\leq j\leq J(i)$, discard $\mathcal{P}_{i}$ if any of the following cases holds.
  1. 1.
    $P_{i}^{j}$ has a disk sublist for which there exists no disk in $\mathcal{S}_{j-1}\setminus\mathcal{G}_{j-1}$ whose boundary contains all endpoints in the disk sublist.
  2. 2.
    The list of endpoints in $P_{i}^{j}$ is not one of those in proposition 2.
  3. 3.
    $P_{i}^{j}$ is an upper (resp. lower) cylinder, its fragments with endpoints on only one end of $P_{i}^{j}$ and with at least one upper endpoint are at higher level than the fragments of $P_{i}^{j}$ they conflict with that have endpoints on both ends of $P_{i}^{j}$ and at least one upper endpoint. (lemma 3)
  4. 4.
    $P_{i}^{j}$ is an upper (resp. lower) cylinder with a fragment with endpoints on both ends of $P_{i}^{j}$, two halftwists with respect to $P_{i}^{j}$ and at least one upper endpoint which is at higher level than the other fragments with endpoints on both ends of $P_{i}^{j}$ and at least one upper endpoint. (lemma 3)
  5. 5.
    $P_{i}^{j}$ is an upper (resp. lower) cylinder, its fragments with endpoints on only one end of $P_{i}^{j}$ and with at least one lower endpoint are at lower level than the fragments of $P_{i}^{j}$ they conflict with that have endpoints on both ends of $P_{i}^{j}$ and at least one lower endpoint. (lemma 3)
  6. 6.
    $P_{i}^{j}$ is an upper (resp. lower) cylinder with a fragment with endpoints on both ends of $P_{i}^{j}$, two halftwists with respect to $P_{i}^{j}$ and at least one lower endpoint which is at lower level than the other fragments with endpoints on both ends of $P_{i}^{j}$ and at least one lower endpoint. (lemma 3)
  7. 7.
    There exists a cylinder $P_{i}^{k}$ in $\mathcal{P}_{i}$ that has the same label as $P_{i}^{j}$. And for a pair of fragments $F_{s}$ at level $s$ and $F_{t}$ at level $t>s$ in $P_{i}^{j}$ that are conflicting with respect to the spine, there exists in $\ell$ an upper (resp. lower) endpoint $x$ at level lower (resp. higher) than $s$ of a fragment $g\in P_{i}^{k}$ that lies between two upper (resp. lower) endpoints of $F_{s}$ and/or $F_{t}$ that belong to the same spine sublist of $P_{i}^{j}$. (proposition 3, corollary 3)
  8. 8.
    Tracing the disks which the fragments of $P_{i}^{j}$ have endpoints on using 1 reveals that the endpoints of all fragments of $P_{i}^{j}$ lie on a single disk $D$ and no pair of fragments of $P_{i}^{j}$ conflicts with respect to the boundary of $D$. (lemma 4)
  9. 9.
    There exists a cylinder $P_{i}^{h}$ with $h<j$ in $\mathcal{P}_{i}$, such that one of the following cases holds:
    • Both $P_{i}^{h}$ and $P_{i}^{j}$ are upper (resp. lower) cylinders, and some lower (resp. upper) endpoints of $P_{i}^{h}$ interact with some upper (resp. lower) endpoints of $P_{i}^{j}$, and the fragments of $P_{i}^{h}$ interacting are at lower (resp. upper) level than the interacting fragments of $P_{i}^{j}$ they conflict with. (proposition 4)
    • $P_{i}^{h}$ is a lower (resp. upper) cylinder, and $P_{i}^{j}$ is an upper (resp. lower) one, and there exists a fragment of $P_{i}^{j}$ which has an upper (resp. lower) endpoint in between two upper (resp. lower) endpoints of fragments of $P_{i}^{h}$ belonging to the same spine sublist of $P_{i}^{h}$. (proposition 5)
    • $P_{i}^{h}$ is an upper (resp. lower) cylinder, and $P_{i}^{j}$ is a lower (resp. upper) one, and there exists a fragment of $P_{i}^{h}$ which has an upper (resp. lower) endpoint in between two upper (resp. lower) endpoints of fragments $g_{1},g_{2}$ of $P_{i}^{j}$ belonging to the same spine sublist of $P_{i}^{j}$ such that there is no path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk with endpoints of $g_{1},g_{2}$ on its boundary which only crosses over (resp. under) $P_{i}^{h}$ or below (resp. above) the fragments embedded in $\mathcal{S}_{0}$ [^1] To detect if a path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk $D$ with endpoints of $g_{1},g_{2}$ on its boundary only crosses over a cylinder $P_{i}^{h}$, it is enough to trace the boundary of $D$ with 1 and check if there exists a path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ which includes at least one fragment embedded on a cylinder $P_{i}^{k}$ with $k>h$ and no fragments embedded on $\mathcal{S}_{l}$ for $l<i$ which conflict with $f$. (proposition 5)
    • $P_{i}^{h}$ is an upper (resp. lower) cylinder with a fragment with an upper (resp. lower) endpoint interacting with a fragment with a lower (resp. upper) endpoint of $P_{i}^{j}$ and $P_{i}^{j}$ is a lower (resp. upper) cylinder. (corollary 4)
If $\mathcal{P}_{i}$ has not been discarded, embed the fragments in $P_{i}^{j}$ in a cylinder $P_{i}^{j}$ that is upper or lower according to the label given to $P_{i}^{j}$. Glue $P_{i}^{j}$ to the disk(s) of $\mathcal{S}_{j-1}\setminus\mathcal{G}_{j-1}$ which were traced in case 4, obtaining the surface $\mathcal{S}_{j}.$ Call $\mathcal{G}_{j}$ the subgraph of $\mathcal{G}$ induced by $\mathcal{G}_{j-1}$ and the fragments in $P_{i}^{j}$. Consider the set $P_{i}^{j+1}\in\mathcal{P}_{i}$ and repeat the process.
If $P_{i}^{j+1}$ does not exist, then the algorithm ends successfully.
If $\mathcal{P}_{i}$ is discarded, consider $\mathcal{P}_{i+1}$. If $\mathcal{P}_{i+1}$ does not exist, set $c=c+1$. If $c=n-1$, the algorithm ends unsuccessfully.
Example 3.
Consider three fragments $A,B,C$, that is $N=3$. Then there are $\text{e}^{-1}\sum_{m=0}^{\infty}\frac{m^{3}}{m!}=5$ partitions $P_{1}=\{\{A\},\{B\},\{C\}\}$, $P_{2}=\{\{A\},\{B,C\}\}$, $P_{3}=\{\{B\},\{A,C\}\}$, $P_{4}=\{\{C\},\{A,B\}\}$, $P_{5}=\{\{A,B,C\}\}$. That is, $J(1)=3$, $J(2)=J(3)=J(4)=2$, $J(5)=1$.
Some examples of cylinders are $P_{1}^{3}=\{C\}$, $P_{2}^{1}=\{A\}$, $P_{2}^{2}=\{B,C\}$, $P_{3}^{1}=\{B\}$.
There are $42^{J(2)}=2^{2}=$ 4 possibilities to label the cylinders of $P_{2}$ as upper or lower cylinders: $\{\{A\}^{U},\{B,C\}^{U}\}$, $\{\{A\}^{U},\{B,C\}^{L}\}$, $\{\{A\}^{L},\{B,C\}^{U}\}$, $\{\{A\}^{L},\{B,C\}^{L}\}$.
Example 4.
If a cylinder contains a single edge $P_{i}^{j}=\{A\}$ on level $b$ the spine sublist of the fragments in $P_{i}^{j}$ is $(A_{b}^{1},A_{b}^{1})$. Since $P_{i}^{j}$ contains only two endpoints, there is only one possibility of dividing the spine sublist into disk sublists, namely $(A_{b}^{1}/A_{b}^{1})$. That leaves eight combinations of labelling the endpoints as well as the cylinder, of which four are labeled as upper cylinders $P_{i}^{j,U}$ and four as lower cylinders $P_{i}^{j,L}$.
That is, if the leveled graph has only three levels and there is only one fragment on level $b=3$, for $c=1$ the algorithm steps through the following partitions:
$\mathcal{P}_{1}$ =$=\{(A_{3,u}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{2}$ =$=\{(A_{3,u}^{1}/A_{3,l}^{1})^{U}\},$ $\mathcal{P}_{3}$ =$=\{(A_{3,l}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{4}$ =$=\{(A_{3,l}^{1}/A_{3,l}^{1})^{U}\},$
$\mathcal{P}_{5}$ =$=\{(A_{3,u}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{6}$ =$=\{(A_{3,u}^{1}/A_{3,l}^{1})^{L}\},$ $\mathcal{P}_{7}$ =$=\{(A_{3,l}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{8}$ =$=\{(A_{3,l}^{1}/A_{3,l}^{1})^{L}\}.$
The following theorem is a consequence of the results of the previous sections.
Theorem 1.
Let $\mathcal{G}$ be a hamiltonian leveled spatial graph. If algorithm 1 applied to $\mathcal{G}$ ends successfully, its output is an oriented closed surface $\mathcal{S}\subset\mathbb{R}^{3}$ without self-intersections where $\mathcal{G}$ is cellular embedded.
Algorithm 1 requires the spine to be a hamiltonian cycle. However, we can generalize the algorithm to general (non-hamiltonian) leveled embeddings.
If we assume that the spine of the leveled spatial graph is a non-hamiltonian cycle, fragments do not need to be edges, but they all are planar connected graphs by definition. In order to place handles under such spatial fragments, we need to attach $n$-punctured spheres instead of $2$-punctured spheres (i.e. cylinders) to build the surface. Note that an $n$-punctured sphere can be obtained by gluing $n-1$ cylinders together.
Algorithm 2.
Let $\mathcal{G}$ be a leveled spatial graph with spine $C$. Let $e$ be any edge of $C$ and choose a spanning tree $T$ of $\mathcal{G}$ which contains $C-e$. The choice of $T-(C-e)$ is independent of the choice of $e$.
Embed $T\cup e$ in a sphere $\mathcal{S}_{0}$, such that $C$ is the equator of $\mathcal{S}_{0}$. This embedding is cellular, because $T\cup e$ is a connected trivial spatial graph. Contracting all edges in $T-C$ gives a hamiltonian leveled spatial graph $\mathcal{G}^{\prime}$ to which algorithm 1 can be applied. If the algorithm terminates successfully, a cellular embedding $(\mathcal{G}^{\prime},\mathcal{S}^{\prime})$ is obtained. The edge contraction can be reversed by vertex splitting and subdividing edges in $\mathcal{S}^{\prime}$ to obtain a cellular embedding $(\mathcal{G},\mathcal{S})$.
The following theorem is a consequence of theorem 1 and algorithm 2.
Theorem 2.
Let $\mathcal{G}$ be a leveled spatial graph. If algorithm 2 applied to $\mathcal{G}$ ends successfully, its output is an oriented closed surface $\mathcal{S}\subset\mathbb{R}^{3}$ without self-intersections where $\mathcal{G}$ is cellular embedded.
Note that the underlying abstract graph of a leveled embedding consists of at most one non-planar 2-connected component, by definition. For example, the graph $K_{5}\_K_{5}$, obtained by joining two copies of $K_{5}$ by an edge (see fig. 11), does not admit a leveled embedding, because it contains two non-planar 2-connected components. However, if multiple leveled embeddings are attached to each other via trivial spatial graphs, the algorithms presented before can still be applied. We call this generalization of leveled embeddings multi-leveled embeddings:
Definition 6.
A connected spatial graph $\mathcal{G}$ that can be constructed as described in the following is multi-leveled.
Consider a tree $T$ and the bipartite partition $X_{1},X_{2}$ of the vertices of $T$. To every vertex $v_{1}^{i}\in X_{1}$ assign a leveled spatial graph $\mathcal{G}_{i}$, $1\leq i\leq|X_{1}|$. The $\mathcal{G}_{i}$ are called blocks of the embedding. To every vertex $v_{2}^{j}\in X_{2}$, $1\leq j\leq|X_{2}|$, assign a connected planar graph $H_{j}$ with at least $\deg(v_{2}^{j})$ vertices. For each vertex $v_{1}^{i}$ adjacent to $v_{2}^{j}$, identify some vertices of $H_{j}$ with some vertices of the block $\mathcal{G}_{i}$ such that identified vertices belong to exactly one planar graph and one block. (See fig. 10)
Figure 10: A multi-leveled embedding with three blocks and two connected planar graphs. On the top left, the tree associated to the embedding.
Figure 11: A multi-leveled embedding of the graph $K_{5}\_K_{5}$ with two blocks, each isomorphic to $K_{5}$, and one planar graph, isomorphic to $K_{2}$.
Theorem 3.
Let $\mathcal{G}$ be a spatial graph with a multi-leveled embedding with blocks $\mathcal{G}_{1},\dots,\mathcal{G}_{n}$. If:
  1. (1)
    algorithm 2 applied to $\mathcal{G}_{i}$ produces a surface $\mathcal{S}_{i}$ for each $1\leq i\leq n$,
  2. (2)
    every $H_{j}$ can be cellular embedded in a $d_{j}$-punctured sphere $S_{d_{j}}$ such that the vertices of $H_{j}$ identified with vertices $w_{j}^{1},\dots,w_{j}^{t}$ of block $\mathcal{G}_{i}$ all lie in the boundary of a puncture of $S_{d_{j}}$,
  3. (3)
    $w_{j}^{1},\dots,w_{j}^{t}$ all lie on the boundary of a single disk $D_{i,j}$ of $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$,
  4. (4)
    the disks can be chosen such that $D_{i,j}\neq D_{i^{\prime},j^{\prime}}$ for $i,j\neq i^{\prime},j^{\prime}$,
then there exists a closed oriented surface $\mathcal{S}$ such that $(\mathcal{G},\mathcal{S})$ is cellular.
Proof.
The surface $\mathcal{S}$ is constructed by attaching the punctured sphere $S_{d_{j}}$ to $\mathcal{S}_{i}$. Therefore the boundary of the puncture where vertices of $H_{j}$ are identified with vertices of $\mathcal{G}_{i}$ and the boundary of the cell $D_{i,j}$ are identified and the interior of $D_{i,j}$ removed. The embeddings of $H_{j}$ in $S_{d_{j}}$ and of $\mathcal{G}_{i}$ on $\mathcal{S}_{i}$ are cellular by assumption. Moreover, condition $(3)$ together with lemma 4 ensure that the embedding $(\mathcal{G},\mathcal{S})$ is also cellular. ∎
4.3.1 Results of the algorithm
The algorithm gives a sufficient condition for a leveled spatial graph to be cellular embedded in an oriented surface. To investigate in how far the successful termination of the algorithm is also a necessary condition, we tried to construct a leveled spatial graph $\mathcal{G}$ for which the algorithm ends unsuccessfully for all possible choices of the spine. We have not been able to construct such an example. In all graphs we studied, it was always possible to find a spine for which theorem 1, theorem 2, proposition 1, or corollary 2 apply. As an example, consider the spatial graph $\mathcal{G}$ described in example 1. Although for the given spine the algorithm does not find a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds, it is possible to choose a non-hamiltonian spine for $\mathcal{G}$, shown in fig. 12, for which algorithm 2 terminates successfully. Furthermore, a different hamiltonian spine for $\mathcal{G}$ can be chosen, shown in fig. 13, which yields a leveled embedding with five levels. Corollary 2 guarantees the existence of a surface where $\mathcal{G}$ cellular embeds, and indeed algorithm 1 terminates successfully if applied to this leveled embedding. Note that it follows from this example that all hamiltonian leveled spatial graphs with a spine list of the form $(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},\dots,n^{1},n^{2},n^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},\dots,n^{3},n^{2},n^{1})$ for $n\in\mathbb{N}$ can be cellular embedded by applying algorithm 1 to a different choice of the spine, analogously to the example.
Figure 12: Left: the new spine chosen for $\mathcal{G}$ in red. Right: the leveled embedding with the new spine.
Figure 13: On the left in red, the new hamiltonian spine chosen for $\mathcal{G}$. On the right, the leveled embedding with the new spine.
It is important to highlight that admitting a leveled embedding for a spatial graph $\mathcal{G}$ is not a necessary condition for the existence of a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds. Indeed, the trefoil knot with an unknotting tunnel is an example of a spatial graph that does not admit a leveled embedding, because for every unknotted cycle chosen as spine, the fragment will cross the spine non-trivially. It does though have a cellular embedding in a torus, as shown in fig. 14.
Figure 14: A trefoil knot with unknotting tunnel cellular embeds in a torus but does not admit a leveled embedding.
In [1], it is shown that any spatial graph $\mathcal{G}$ can be embedded in a 3-book, where the vertices lie on the back of the 3-book and edges may be contained in several leaves (i.e. the pages of the 3-book). Placing a vertex on every edge of $\mathcal{G}$ which intersects the back of the 3-book and adding an edge between two consecutive vertices on the back which are not already adjacent yields a leveled embedding of a spatial graph $\mathcal{G}^{\prime}$ in three levels such that $\mathcal{G}\subset\mathcal{G}^{\prime}$. In other words, every spatial graph has a super-graph with a leveled embedding for which algorithm 2 ends successfully.
5 Outlook
For a given spatial graph $\mathcal{G}$, if algorithm 2 ends unsuccessfully for every choice of spine $C$ of a leveled embedding of $\mathcal{G}$, it does not imply that $\mathcal{G}$ cannot be cellular embedded in any surface. However, in every example we studied, it is possible to choose a different spine for which algorithm 2 ends successfully. This motivates the following conjectures.
Conjecture 1.
If a spatial graph $\mathcal{G}$ has a leveled embedding and admits a cellular embedding in some surface, there is a choice of spine for which algorithm 2 ends successfully.
Conjecture 2.
If a spatial graph $\mathcal{G}$ has a leveled embedding, then a cycle can be found that is the spine of a leveled embedding of $\mathcal{G}$ for which algorithm 2 ends successfully.
A similar conjecture can be stated for hamiltonian leveled spatial graphs. The situation shown in fig. 13 is representative for all examples of hamiltonian leveled spatial graphs that we studied.
Conjecture 3.
If a spatial graph $\mathcal{G}$ has a hamiltonian leveled embedding, then there exists a choice of hamiltonian spine and corresponding hamiltonian leveled embedding of $\mathcal{G}$ for which algorithm 1 ends successfully.
The algorithms 1 and 2, while giving a constructive procedure to find a surface $\mathcal{S}$ where a spatial graphs cellular embeds, are intractable to actually implement for spatial graphs with a high number of levels. For a practical implementation, some heuristics would be needed to optimize the algorithm, based for example on which choice of cylinders is more likely for the algorithm to end successfully.
Cellular embeddings of graphs are related to circular embeddings. A circular embedding is a cellular embedding $(\mathcal{G},\mathcal{S})$ where the boundaries of the faces of $\mathcal{S}\setminus\mathcal{G}$ are cycles of $\mathcal{G}$. A famous open conjecture in graph theory states that every 2-connected graph has a circular embedding [6]. In case of a 2-connected hamiltonian leveled spatial graph $\mathcal{G}$ for which algorithm 1 ends successfully, it is possible to determine whether the resulting cellular embedding is circular by applying 1 to all faces of the cellular embedding: if every edge of $\mathcal{G}$ appears in the boundary of exactly two faces, then the embedding is circular. The embedding $(\mathcal{G},\mathcal{S})$ is ensured to be circular if no cylinder attachments merge disks in the construction of $\mathcal{S}$ as performed in algorithm 1.
Investigating which embeddings obtained using algorithm 2 are circular would give insights in the circular embedding conjecture for 2-connected graphs with a multi-leveled embedding.
6 Acknowledgement
The authors want to thank Joanna A. Ellis-Monaghan for discussions that motivated this research.

### v2

@@
 Abstract
-Finding a closed orientable surface $\mathcal{S}$ embedded in $\mathbb{R}^{3}$ where a given spatial graph $\mathcal{G}\subset\mathbb{R}^{3}$ cellular embeds is in general not possible.
-We therefore restrict our interest to the special class of spatial graphs that are leveled.
-We show that for leveled spatial graphs with a small number of levels, a surface $\mathcal{S}$ can always be found.
-The argument is based on the idea of decomposing $\mathcal{G}$ into subgraphs that can be placed on a sphere and on handles that are attached to the sphere, together forming an embedding of $\mathcal{G}$ in $\mathcal{S}$.
+For a given spatial graph $\mathcal{G}\subset\mathbb{R}^{3}$, we would like to find a closed orientable surface $\mathcal{S}$ embedded in $\mathbb{R}^{3}$ in which $\mathcal{G}$ is cellular embedded.
+However, for general $\mathcal{G}$ this is not possible.
+We therefore define a property of spatial graphs, called leveled, to show that for leveled spatial graphs with a small number of levels, a surface $\mathcal{S}$ can always be found.
+The argument is based on decomposing $\mathcal{G}$ into spatial subgraphs that can be placed on a sphere and on cylinders attached as handles, in such a way that the resulting surface contains a cellular embedding of $\mathcal{G}$.
 We generalize the procedure to an algorithm that, if successful, constructs $\mathcal{S}$ for leveled spatial graphs with any number of levels.
@@
 Cellular embeddings, as a natural generalization of planar graphs, have in particular been studied in topological graph theory and combinatorics.
-They provide a combinatorial description of graph embeddings in surfaces [9], and they are at the core of long-standing open conjectures such as the circular (or strong) embedding conjecture [6].
+They provide a combinatorial description of graph embeddings in surfaces Sta (78), and they are at the core of long-standing open conjectures such as the circular (or strong) embedding conjecture Jae (85).
 It is possible to cellular embed any abstract graph in a closed orientable surface.
@@
 This motivates the question for which spatial graphs a closed orientable surface embedded in $\mathbb{R}^{3}$ can be found, such that the graph cellular embeds in it, and how the surfaces can be constructed.
-Given that graphs on surfaces have mainly be studies in topological graph theory that, in contrast to knot theory, traditionally considers all possible embeddings of an abstract graph, the scenario involving specific realizations of graphs within three-dimensional space, where surfaces must not intersect, remains underexplored.
+Given that graphs on surfaces have mainly be studied in topological graph theory that, in contrast to knot theory, traditionally considers all possible embeddings of an abstract graph, the scenario involving specific realizations of graphs within three-dimensional space, where surfaces must not intersect, remains underexplored.
 This is despite the setting being most relevant for applications taking place in the 3-dimensional Euclidean world.
-For example, building a graph-like structure using self-assembly of DNA molecules by origami folding requires first finding an unknotted A-trail through the desired spatial graph [4], [7].
-When the graph is eulerian and cellular embedded in a surface, an unknotted A-trail can be found by following the boundary of the disks [4].
+For example, building a graph-like structure using self-assembly of DNA molecules by origami folding requires first finding an unknotted A-trail through the desired spatial graph EMPS+ (17), MAG+ (20).
+When the graph is eulerian and cellular embedded in a surface, an unknotted A-trail can be found by following the boundary of the disks EMPS+ (17).
 Since we do not require the surface to be standardly embedded (e.g.
@@
 However, our goal is to construct a surface for a given spatial graph.
-Note that rotation systems, although equivalent to cellular embeddings for abstract graphs [9], do not ensure the described surface to be free of self-intersections.
+Note that rotation systems, although equivalent to cellular embeddings for abstract graphs Sta (78), do not ensure the described surface to be free of self-intersections.
 A natural approach to construct a surface for a given spatial graph is to glue together non-intersecting properly embedded disks that are bounded by some cycles of the graph.
-A sufficient albeit unnecessarily strong condition for the existence of such disks is $k$-flatness of the spatial graph [8].
+A sufficient albeit unnecessarily strong condition for the existence of such disks is $k$-flatness of the spatial graph NT (04).
 Indeed, it suffices to find a collection of cycles that cover all edges of the graph such that each edge is covered at most twice, and that bound disks satisfying the property above.
@@
 If the two boundary components of the cylinder are attached to the same disk, it must be ensured that the graph also runs through the meridian of the resulting handle.
-Furthermore, conditions need to be found that guarantee that the construction results in an embedded surface, i.e., handles should be attached in such a way that they do not intersect the already constructed surfasce parts.
+Furthermore, conditions need to be found that guarantee that the construction results in an embedded surface, i.e., handles should be attached in such a way that they do not intersect the already constructed surface parts.
 For such a procedure to work, we require the spatial graph to have some properties and we call such spatial graphs leveled.
-They can be viewed as a generalization of planar graphs and are distinct from but related to open book embeddings [2].
+They can be viewed as a generalization of planar graphs and are distinct from but related to open book embeddings BB (24).
 We show in Proposition 1 that all leveled spatial graphs with up to four levels can be cellular embedded in a surface that itself is embedded in $\mathbb{R}^{3}$ by presenting a constructive argument.
@@
 Graphs are allowed to have loops and parallel edges.
-The fragments (or bridges) of a graph $G$ with respect to a cycle $C$ of $G$ are the closures of the connected components of $G-C$ [10].
-Two fragments are conflicting if they either have pairs of endpoints on $C$ that alternate, or have at least three endpoints in common [11].
+The fragments (or bridges) of a graph $G$ with respect to a cycle $C$ of $G$ are the closures of the connected components of GCG-C Tut (56).
+Two fragments are conflicting if they either have pairs of endpoints on $C$ that alternate, or have at least three endpoints in common Wes (96).
 The conflict graph of a graph $G$ with respect to a cycle $C$ is the graph that has a vertex for each fragment of $G$ with respect to $C$, and two vertices are adjacent if and only if the corresponding fragments conflict.
@@
 A diagram of a spatial graph $\mathcal{G}$ is the image of a projection of $\mathcal{G}$ onto $\mathbb{R}^{2}$ such that all intersections in the image are transversal, no vertices and at most two points are mapped to a crossing, and each double point is assigned as an over- or under-crossing.
-We focus on a special class of spatial graphs, the leveled graphs:
+We focus on a special class of spatial graphs, the leveled graphs.
 Definition 1.
@@
 Note that conflicting fragments must be on different levels.
-This definition allows to assign the fragments of a leveled spatial graph to their level. The fragments that do not cross over any other fragments are at level $1$, denoted $L_{1}$.
+Also, it is not possible that two fragments $f,g$ are such that $f>g$ and $g>f$.
+Indeed, if this was the case, the interior $\mathring{D_{f}}$ of $D_{f}$ could not be disjoint from $D_{g}$.
+This definition allows us to assign the fragments of a leveled spatial graph to their level. The fragments that do not cross over any other fragments are at level $1$, denoted $L_{1}$.
 Remark 1.
-If the leveled spatial graph $\mathcal{G}$ has $n$ levels $L_{i},1\leq i\leq n$, there is an action of the cyclic group $C_{n}=\langle x\mid x^{n}=1\rangle$ on the partition of levels =$\mathcal{L}=\{L_{1},\dots,L_{n}\}$.
+If the leveled spatial graph $\mathcal{G}$ has $n$ levels $L_{i},1\leq i\leq n$, there is an action of the cyclic group $C_{n}=\langle x\mid x^{n}=1\rangle$ on the partition of levels $\mathcal{L}=\{L_{1},\dots,L_{n}\}$.
 The action is given by sending $x\cdot L_{i}$ to $L_{i+1(\mathrm{mod}\ n)}.$
-We say that a leveled spatial graph $\mathcal{G}$ is in representative leveling if for each fragment $f$ at level $i>1$, there exists a fragment $g$ at level $i-1$ such that $f$ and $g$ conflict.
+We say that a leveled spatial graph $\mathcal{G}$ is in representative leveling if for each fragment $f$ at level $i>1$, there exists a fragment $g$ at level i1i-1 such that $f$ and $g$ conflict.
 In other words, $\mathcal{G}$ is in representative leveling if every fragment belongs to the lowest possible level.
@@
 Every connected trivial spatial graph is a leveled graph with exactly one level.
-Consequently, all planar graphs admit a leveled embedding, but this is not true for all graphs.
-For example, the graph $K_{5}\_K_{5}$ that is obtained by joining two copies of $K_{5}$ by an edge (fig.
-11) cannot be embedded as a leveled graph, since every choice of a cycle implies a nonplanar fragment.
+Consequently, all planar graphs admit a leveled embedding, but that is not true for all graphs.
+For example, the graph $K_{5}\_K_{5}$ that is obtained by joining two copies of $K_{5}$ by an edge (Figure 13) cannot be embedded as a leveled graph, since every choice of a cycle implies a non-planar fragment.
 Therefore, no spine can be found for $K_{5}\_K_{5}$.
@@
 If the two fragments that have the vertex $v$ in common are non-conflicting, following $C$ starting from $v$ runs through all endpoints of one fragment before running through all endpoints of the other.
-Splitting $v$ as described at above does not introduce an alternation of the endpoints, and therefore keeps the fragments non-conflicting.
-By repeatedly splitting vertices on $C$, their degree can be reduced to three. ∎
+Splitting $v$ as described above does not introduce an alternation of the endpoints, and therefore keeps the fragments non-conflicting.
+∎
+By repeatedly splitting vertices on the spine of a leveled spatial graph in the way described in Lemma 1, their degree can be reduced to three.
 A hamiltonian leveled spatial graph $\mathcal{H}$ is a leveled spatial graph whose spine is a hamiltonian cycle of the underlying abstract graph.
@@
 If several fragments share an endpoint, the order in the spine list is chosen such that the alternation of fragment labels is avoided.
-The spine list is unique up to permutation of the labels of parallel edges and cyclic permutations of its symbols, which corresponds to different choices of start points on the spine, see fig.
-1.
+The spine list is unique up to permutation of the labels of parallel edges and cyclic permutations of its symbols, which corresponds to different choices of start points on the spine, see Figure 1.
 Figure 1: A hamiltonian leveled spatial graph with spine list $(1^{1},2^{2},3^{1},3^{2},2^{1},3^{2},3^{1},1^{2},2^{1},2^{2},2^{3},1^{2},2^{3},1^{1})$.
@@
 Assume $\mathcal{G}$ is in representative leveling.
-Let $L_{i}$ be the union of all fragments of level $i$, and let $L_{1,2}:=C\cup L_{1}\cup L_{2}$ be the subgraph of $\mathcal{G}$ that consists of the spine $C$ of $\mathcal{G}$, the fragments of level $1$ and the fragments of level $2$.
+Let $L_{i}$ be the union of all fragments of level $i$, and let $L_{1,2}:=C\cup L_{1}\cup L_{2}$ be the subgraph of $\mathcal{G}$ that consists of the spine $C$ of $\mathcal{G}$, the fragments of level 11 and the fragments of level $2$.
 As $L_{1,2}$ is trivial by definition, there exists an embedded sphere $\mathcal{S}_{0}$ where $L_{1,2}$ is embedded.
 Since embeddings of connected spatial graphs in a sphere are cellular, the resulting embedding $(L_{1,2},\mathcal{S}_{0})$ is cellular.
-Now handles are attached to the upper hemisphere $\mathcal{S}_{u}$ of $\mathcal{S}_{0}$ to construct a surface on which $L_{1,2}\cup L_{3}$ is embedded in.
+Now handles are attached to the upper hemisphere $\mathcal{S}_{u}$ of $\mathcal{S}_{0}$ to construct a surface where $L_{1,2}\cup L_{3}$ is embedded.
 For that, consider the set of disks $D_{u}$ that are the connected components of $\mathcal{S}_{u}\setminus\mathcal{G}$.
@@
 Denote by $F_{i}$ the set of fragments of $L_{3}$ with endpoints in $Cl_{i}$.
-Since $F_{i}$ is trivial, there is a sphere $S_{F_{i}}$ on which $F_{i}$ embeds in.
+Since $F_{i}$ is trivial, there is a sphere $S_{F_{i}}$ where $F_{i}$ embeds.
 For each set of endpoints $p_{k}$ of $F_{i}$ for which $a(p_{k})=D_{p_{k}}$, remove a disk $P_{p_{k}}$ from $S_{F_{i}}$ such that $p_{k}\in\partial P_{p_{k}}$ for all $k$ and $q\notin P_{p_{k}}$ for any endpoint $q\neq p_{k}$.
-This construction yields an $n_{i}$-punctured sphere $\mathcal{S}_{i}$ on which all fragments of $Cl_{i}$ are embedded in.
+This construction yields an $n_{i}$-punctured sphere $\mathcal{S}_{i}$ where all fragments of $Cl_{i}$ are embedded.
 The embedding of $F_{i}$ in $\mathcal{S}_{i}$ is cellular, because the spatial graph $F_{i}$ together with the boundary cycles of the disks $P_{p_{k}}$ is a connected trivial spatial graph.
 Since the punctured spheres $\mathcal{S}_{i}$ are constructed to follow the non-conflicting fragments of level $3$, it is implied that the $\mathcal{S}_{i}$ are not intersecting.
-Removing the interiors of the disks that are in the image of $a$ from $\mathcal{S}_{u}$ allows to construct an oriented closed non-intersecting surface.
-This is done by identifying their boundaries with the boundaries of the punctures of $\mathcal{S}_{i}$ that respects the induced orientations on the boundaries of the surfaces.
+Removing the interiors of the disks that are in the image of $a$ from $\mathcal{S}_{u}$ allows us to construct an oriented closed non-intersecting surface.
+That is done by identifying their boundaries with the boundaries of the punctures of $\mathcal{S}_{i}$ that respect the induced orientations on the boundaries of the surfaces.
 Analogously, $L_{4}$ is (cellular) embedded in $n_{j}$-punctured punctured spheres $\mathcal{S}_{j}$ that are attached to the lower hemisphere $\mathcal{S}_{l}$.
-The punctured spheres on which $L_{4}$ embeds in do not intersect the surface where $L_{1,2}\cup L_{3}$ is embedded.
-This is due to there being four levels only: each level is attached to only one of the hemispheres and the crossing types between the levels are of only one type, making the attachment of handles to the upper hemisphere and to the lower hemisphere independent of one another.
-The result is an embedding of $\mathcal{G}=L_{1}\cup L_{2}\cup L_{3}\cup L_{4}$ in a handlebody of genus $\sum_{i}(n_{i}-1)+\sum_{j}(n_{j}-1)$.
+The punctured spheres where $L_{4}$ embeds do not intersect the surface where $L_{1,2}\cup L_{3}$ is embedded.
+That is due to there being four levels only: each level is attached to only one of the hemispheres and the crossing types between the levels are of only one type, making the attachment of handles to the upper hemisphere and to the lower hemisphere independent of one another.
+The result is an embedding of $\mathcal{G}=L_{1}\cup L_{2}\cup L_{3}\cup L_{4}$ in a surface of genus $\sum_{i}(n_{i}-1)+\sum_{j}(n_{j}-1)$.
 It remains to be seen that the embedding is cellular.
-The embedding of $\mathcal{G}$ in the resulting handlebody is cellular since each $\mathcal{S}_{k}$ is attached to different disks of $\mathcal{S}_{0}\setminus\mathcal{G}$, and $F_{k}$ is cellular embedded in $\mathcal{S}_{k}$, and $L_{1,2}$ is cellular embedded in $\mathcal{S}_{0}$.
+The embedding of $\mathcal{G}$ in the resulting surface is cellular since each $\mathcal{S}_{k}$ is attached to different disks of $\mathcal{S}_{0}\setminus\mathcal{G}$, and $F_{k}$ is cellular embedded in $\mathcal{S}_{k}$, and $L_{1,2}$ is cellular embedded in $\mathcal{S}_{0}$.
 Corollary 1.
@@
 For degrees greater than three, $a$ is uniquely defined if and only if neither fragments of level 3 and fragments of level 2 nor fragments of level 4 and fragments of level 1 have endpoints in common.
-The restriction to four levels in proposition 1 assures both the cellularity of the resulting embedding and the absence of self-intersections of the resulting surface.
+The restriction to four levels in Proposition 1 assures both the absence of self-intersections of the resulting surface and the cellularity of the resulting embedding.
 The construction of the proof of Proposition 1 ensures neither property for a leveled embedding with five or more levels, as can be seen in the following example.
 Example 1.
-Consider the leveled embedding $\mathcal{G}$ shown left in fig. 13, that is given by the spine list
-$(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},3^{1},3^{2},3^{3},4^{1},4^{2},4^{3},5^{1},5^{2},5^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},3^{3},3^{2},3^{1},4^{3},4^{2},4^{1},5^{3},5^{2},5^{1}).$
-We can apply the algorithm in the proof of proposition 1 to embed any four levels in a surface $\mathcal{S}$; without loss of generality choose the levels 1, 2, 4, 5.
+Consider the leveled embedding $\mathcal{G}$ shown left in Figure 15, that is given by the spine list $(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},3^{1},3^{2},3^{3},$ $4^{1},4^{2},4^{3},5^{1},5^{2},5^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},$ $3^{3},3^{2},3^{1},4^{3},4^{2},4^{1},5^{3},5^{2},5^{1}).$ We can apply the algorithm in the proof of Proposition 1 to embed any four levels in a surface $\mathcal{S}$; without loss of generality choose the levels 1, 2, 4, 5.
 This leaves level 3 to be embedded.
@@
 As the three fragments do not conflict in $\mathcal{G}$ with respect to the spine, and $P$ is glued to $D_{u}$ on one end and to $D_{l}$ on the other, the two ends of $P$ are oriented with the same orientation as the spine.
-This is, if the orientation on one boundary component of the cylinder is clockwise, it is counterclockwise for the other.
+That is, if the orientation on one boundary component of the cylinder is clockwise, it is counterclockwise for the other.
 The boundary orientation does therefore not correspond to the orientation that is obtained when considering the two ends of the cylinder as naturally identified.
@@
 Reversing the order of the endpoints in one of the spine sublist results in the disk sublists of $P_{i}$.
-This is, to obtain the disk sublists, the boundary components of the cylinder are considered as naturally identified.
+That is, to obtain the disk sublists, the boundary components of the cylinder are considered as naturally identified.
+Definition 4.
+Let $f$ be a fragment embedded in a cylinder $P$ such that the endpoints of $f$ lie on both ends of $P$.
+Consider the rectangle representation of $P$ given by $0,1]\times[0,2\pi]/\sim$, where the point $(t,0)$ is identified with the point $(t,2\pi)$ for $t\in[0,1].$ We say that $f$ has a right (resp.
+left) halftwist with respect to $P$ if $f$ is ambient isotopic with fixed endpoints to the line joining points $(0,0)$ and $(1,\pi)$ (resp.
+the line joining points $(0,2\pi)$ and $(1,\pi)$).
+(Figure 2)
+Figure 2: A fragment $f$ with a right halftwist with respect to the cylinder $P$.
 Remark 2.
@@
 Therefore, the difference of halftwists with respect to $P_{i}$ of the two fragments is at most 1.
-This is impossible due to the previous argument.
+That is impossible due to the previous argument.
 ∎
@@
   3. 3.
-    $(s_{1},\dots,s_{n},\alpha_{1},\dots\alpha_{k},\beta_{1},\dots,\beta_{h},s_{n},\dots,s_{1}/t_{1},\dots,t_{q},\beta_{1},\dots,\beta_{h},\alpha_{1},\dots,\alpha_{k},t_{q},\dots,t_{1}),$
+    $(s_{1},\dots,s_{n},\alpha_{1},\dots\alpha_{k},\beta_{1},\dots,\beta_{h},s_{n},\dots,s_{1}/t_{1},\dots,t_{q},\beta_{1},\dots,\beta_{h},$
+    $\alpha_{1},\dots,\alpha_{k},t_{q},\dots,t_{1}),$
   4. 4.
-    $(a_{1},\dots,a_{m},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p},\beta_{1},\dots,\beta_{h},a_{m},\dots,a_{1}/\beta_{1},\dots,\beta_{h},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p});$
+    $(a_{1},\dots,a_{m},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p},\beta_{1},\dots,\beta_{h},a_{m},\dots,a_{1}/\beta_{1},\dots,\beta_{h},$
+    $\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p});$
   5. 5.
@@
 The endpoints of the fragments embedded in $P_{i}$ can be moved along the connected component of the boundary of $P_{i}$ to which the endpoints belong to while maintaining their embedding in $P_{i}$.
-Lemma 2 gives conditions for the embedding to be leveled: in order to obtain the lists presented, consider the cylinder $P_{i}$ embedded in $\mathbb{R}^{3}$ but not glued to $\mathcal{S}_{i}$ depicted in fig.
-2.
+Lemma 2 gives conditions for the embedding to be leveled: in order to obtain the lists presented, consider the cylinder $P_{i}$ embedded in $\mathbb{R}^{3}$ but not glued to $\mathcal{S}_{i}$ depicted in Figure 3.
 Shifting the endpoints of the fragments along the boundary components of $P_{i}$ while respecting the conditions given by Lemma 2 yields the six presented lists.
@@
 Lemma 2 and Proposition 2 restrict how fragments are embedded in a cylinder $P_{i}$ and the possible disk sublists of $P_{i}$ in a consecutively embedded leveled spatial graph.
-Proposition 2 allows to extend proposition 1:
+Example 2.
+Consider the cylinder $P$ on the right of Figure 3.
+If we name $a_{1}$ the red fragment with endpoints attached to one end of $P$, $\alpha$ the red fragment with endpoints attached to both ends of $P$, $b_{1},b_{2},b_{3}$ the green fragments with endpoints on both ends of $P$ and $x_{1},x_{2},x_{3}$ the green fragments with endpoints on one end of $P$, then the disk sublists of $P$ would be
+$(a_{1},\alpha,b_{1},b_{2},x_{1},x_{1},b_{3},a_{1}/\alpha,x_{2},x_{2},b_{1},b_{2},b_{3},x_{3},x_{3}),$
+which correspond to the sublists of type $4$ of Proposition 2.
+Proposition 2 allows us to extend Proposition 1:
 Corollary 2.
@@
 $N_{l}$).
-(fig.
-3)
+(Figure 4)
 Figure 4: An upper cylinder $P$ with a fragment $f$ with upper endpoints and a lower cylinder $P^{\prime}$ with a fragment $g$ with upper endpoints.
@@
 Lemma 3.
-Consider a cylinder $P_{i}$ of the consecutively embedding of $\mathcal{G}$ in $\mathcal{S}.$
+Consider a cylinder $P_{i}$ of the consecutively embedded spatial graph $\mathcal{G}$ in $\mathcal{S}.$
   1. 1.
@@
 The other cases are similar.
-We prove the counterpositive: if $x$ is embedded in an upper regular cylinder, then $x$ is at higher level than $a$.
-(fig.
-5)
+We prove the counterpositive: if $x$ is embedded in an upper cylinder, then $x$ is at higher level than $a$.
+(Figure 6)
 If $x$ is embedded in $P$, then it must be at higher level than $a$, by Lemma 3.
@@
 We only need to prove 2, because 1 follows from Proposition 3 and 3 follows from 2 by applying a reflection with respect to the plane where the spine lies.
-We prove the counterpositive: if $P$ and $P^{\prime}$ are upper cylinders and the endpoints of $a,b,x$ are lower endpoints, then $x$ is at lower level than $a$ (fig.
-6).
+We prove the counterpositive: if $P$ and $P^{\prime}$ are upper cylinders and the endpoints of $a,b,x$ are lower endpoints, then $x$ is at lower level than $a$ (Figure 7).
 Let $P$ be an upper cylinder and $x$ have a lower endpoint in between two lower endpoints of $a$ and $b$ in the same disk sublist.
 If $x$ can also be embedded in $P$, then $x$ is at lower level than $b$ due to Lemma 3.
-If $x$ is embedded in another upper consonant cylinder $P^{\prime}$, then it conflicts with either both or exactly one of $a$ and $b$.
-In the latter case, the crossing number of $x$ with the fragment it does not conflict with must be 0.
+If $x$ is embedded in another upper cylinder $P^{\prime}$, then it conflicts with either both or exactly one of $a$ and $b$.
+In the latter case, the crossing number of $x$ with the fragment it does not conflict with must be 0, where the crossing number is the minimum number of crossing points between two fragments up to ambient isotopy.
 If $x$ conflicts with $a$, then $x$ is at lower level than $a$, because else $x$ would need to pierce the cylinder $P$ or cross over $P$, making it cross also fragment $b$; contradicting $x$ having crossing number 0 with $b$.
 If $x$ conflicts only with $b$, it is not possible that $x$ is at higher level than $b$, because then $P^{\prime}$ would cross over $P$, which implies that $x$ crosses over $a$ non-trivially.
-This is a contradiction of the fact that $x$ has crossing number 0 with $a$.
+That is a contradiction of the fact that $x$ has crossing number 0 with $a$.
 The case where $x$ conflicts with both $a$ and $b$ implies that $x$ is at lower level than $a$.
@@
 If $x$ is at higher level than $b$, then $P^{\prime}$ must cross over $P$.
-But this is not possible because $x$ has a lower endpoint in between two lower endpoints of $a$ and $b$.
+But that is not possible because $x$ has a lower endpoint in between two lower endpoints of $a$ and $b$.
 Therefore, $x$ needs first to cross under $a$ or $b$ before crossing over $b$, which is a contradiction to being leveled.
@@
 The ends of the two cylinders where $e$ and $e^{\prime}$ lie do not intersect each other as the plane where the spine lies separates them.
-(fig.
-7)
+(Figure 8)
 We cover cases 2, 3 and 4 in Proposition 4, Proposition 5 and Corollary 4 respectively.
@@
 If the fragments $f_{1},\dots,f_{s}$ of $P$ which conflict with a fragment of $g_{1},\dots,g_{t}$ of $P^{\prime}$ are at lower level than the fragments of $P^{\prime}$ they conflict with, then the lower endpoints of $f_{1},\dots,f_{s}$ do not interact with the upper endpoints of $g_{1},\dots,g_{t}$ of $P^{\prime}$ they conflict with.
-(See fig.
-8 for the counterpositive.)
+(See Figure 9 for the counterpositive.)
 Proof.
@@
 In this section, we investigate which hamiltonian leveled spatial graphs that are consecutively embedded are cellular.
+Recall that the subgraph of a consecutively embedded leveled spatial graph $\mathcal{G}$ embedded in a surface $S_{i}$, built successively by gluing the ends of cylinders to a sphere, is called $\mathcal{G}_{i}$.
 Lemma 4.
@@
   1. (1)
-    If $P_{i}$ is glued to two different disks $D_{1},D_{2}$ of the cellular embedding of $\mathcal{S}_{i-1}\setminus\mathcal{G}_{i-1}$, then the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular.
+    If $P_{i}$ is glued to two different disks $D_{1},D_{2}$ of the cellular embedding of $\mathcal{S}_{i-1}\penalty 10000\ \setminus\penalty 10000\ \mathcal{G}_{i-1}$, then the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular.
   2. (2)
-    If $P_{i}$ is glued to a single disk $D$, then the resulting embedding is cellular if and only if there exists at least one conflicting pair of fragments in $f_{1},\dots f_{t}$ in the conflict graph with respect to the boundary of $D.$
+    If $P_{i}$ is glued to a single disk $D$, then the resulting embedding is cellular if and only if there exists at least one conflicting pair of fragments in $f_{1},\dots,f_{t}$ in the conflict graph with respect to the boundary of $D.$
 Proof.
@@
 In our case, loops are fragments and two fragments can cross each other only in $t_{0}$.
-Therefore, $((T,t_{o}),P_{i})$ is cellular if and only if there exist at least two fragments $f_{1},f_{2}$ such that their endpoints alternate on the boundary of $D.$
+Therefore, the embedding of $P_{i}$ in the torus $T$ with marked point $t_{o}$ is cellular if and only if there exist at least two fragments $f_{1},f_{2}$ such that their endpoints alternate on the boundary of $D.$
 Example 3.
 Consider the leveled spatial graph $\mathcal{G}$ given by $(3,1,4,2,1,3,4,2).$ We would like to embed $\mathcal{G}$ consecutively such that all endpoints are upper endpoints, fragment $2$ is embedded in the upper hemisphere of a sphere $\mathcal{S}_{0}$, fragment $1$ is embedded in a lower cylinder $P_{1}$ glued to the two disks formed by fragment $2$ (notice that this merges the two disks of the upper hemisphere into a single disk $D$) and fragments $3$ and $4$ are embedded together in an upper cylinder $P_{2}$ glued to $D$.
-See fig.
-9 for a picture of the embedding.
+See Figure 10 for a picture of the embedding.
 This embedding of $\mathcal{G}$ in $\mathcal{S}_{2}$ is not cellular, because fragments $3$ and $4$ do not conflict with respect to the boundary of $D$ by Lemma 4.
@@
 Suppose that $\mathcal{G}_{P_{i}}$ is cellular embedded in $P_{i}$ and $t$ is the number of fragments of $\mathcal{G}_{P_{i}}\subset\mathcal{G}_{i}$.
-Then, $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $n-2+t$ disks.
+Then, $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has n2+tn-2+t disks.
 Proof.
 The cellular embedding of $\mathcal{G}_{P_{i}}$ in $P_{i}$ has $t$ disks.
-Indeed, the first fragment of $\mathcal{G}_{P_{i}}$ which is attached to both ends of $P_{i}$ cuts an annulus into a disk, and every other fragment of $\mathcal{G}_{P_{i}}$ cuts an existing disk in two disks, adding one disk to the total per fragment.
-If $P_{i}$ is glued to two different disks of $\mathcal{S}_{i-1}$, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $n-2+t$ disks.
-If $P_{i}$ is glued to a single disk $D$ and the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular as seen in lemma 4, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $(n-1)+(t-1)$ disks.
+Indeed, the first fragment of $\mathcal{G}_{P_{i}}$ which is attached to both ends of $P_{i}$ transforms an annulus into a disk, and every other fragment of $\mathcal{G}_{P_{i}}$ transforms an existing disk in two disks, adding one disk to the total per fragment.
+If $P_{i}$ is glued to two different disks of $\mathcal{S}_{i-1}$, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has n2+tn-2+t disks.
+If $P_{i}$ is glued to a single disk $D$ and the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular as seen in Lemma 4, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has (n1)+(t1)(n-1)+(t-1) disks.
 Indeed, for each end of $P_{i}$, the disk of $P_{i}\setminus\mathcal{G}_{P_{i}}$ that contains $\partial P_{i}$ which is not part of the spine is merged with $D$.
@@
 ∎
-Given that the embedding is cellular, we adapt the “Face tracing algorithm” of a cellular graph embedding explained in [5] to be able to trace the boundaries of its disks.
+Given that the embedding is cellular, we adapt the “Face tracing algorithm” of a cellular graph embedding explained in GT (01) to be able to trace the boundaries of its disks.
 Fact 1 (Face tracing algorithm).
@@
 Let $\mathcal{F}_{-\{c,c+1\}}$ be the set of $N$ fragments of $\mathcal{G}$ that are not in level $c$ and $c+1$.
-Generate all possible partitions $P_{k},1\leq k\leq\text{e}^{-1}\sum_{m=0}^{\infty}\frac{m^{N}}{m!}$ (Dobińsky’s formula [3]), of $\mathcal{F}_{-\{c,c+1\}}$ into disjoint nonempty subsets.
+Generate all possible partitions $P_{k},1\leq k\leq\text{e}^{-1}\sum_{m=0}^{\infty}\frac{m^{N}}{m!}$ (Dobińsky’s formula Dob (77)), of $\mathcal{F}_{-\{c,c+1\}}$ into disjoint nonempty subsets.
 That is, the elements of $P_{k}$ are subsets $P_{k}^{j}$, $1\leq j\leq J(k)$ of fragments, where $J(k)$ is the number of the subsets in $P_{k}$.
@@
 under) $P_{i}^{h}$ or below (resp.
-above) the fragments embedded in $\mathcal{S}_{0}$ [^1] To detect if a path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk $D$ with endpoints of $g_{1},g_{2}$ on its boundary only crosses over a cylinder $P_{i}^{h}$, it is enough to trace the boundary of $D$ with 1 and check if there exists a path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ which includes at least one fragment embedded on a cylinder $P_{i}^{k}$ with $k>h$ and no fragments embedded on $\mathcal{S}_{l}$ for $l<i$ which conflict with $f$.
+above) the fragments embedded in $\mathcal{S}_{0}$.
+[^1] To detect if a path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk $D$ with endpoints of $g_{1},g_{2}$ on its boundary only crosses over a cylinder $P_{i}^{h}$, it is enough to trace the boundary of $D$ with 1 and check if there exists a path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ which includes at least one fragment embedded on a cylinder $P_{i}^{k}$ with $k>h$ and no fragments embedded on $\mathcal{S}_{l}$ for $l<i$ which conflict with $f$.
 (Proposition 5)
@@
 If $P_{i}^{j+1}$ does not exist, then the algorithm ends successfully.
-If $\mathcal{P}_{i}$ is discarded, consider $\mathcal{P}_{i+1}$. If $\mathcal{P}_{i+1}$ does not exist, set $c=c+1$. If $c=n-1$, the algorithm ends unsuccessfully.
+If $\mathcal{P}_{i}$ is discarded, consider $\mathcal{P}_{i+1}$. If $\mathcal{P}_{i+1}$ does not exist, set $c=c+1$. If c=n1c=n-1, the algorithm ends unsuccessfully.
 Example 4.
@@
 Some examples of cylinders are $P_{1}^{3}=\{C\}$, $P_{2}^{1}=\{A\}$, $P_{2}^{2}=\{B,C\}$, $P_{3}^{1}=\{B\}$.
-There are $42^{J(2)}=2^{2}=$ 4 possibilities to label the cylinders of $P_{2}$ as upper or lower cylinders: $\{\{A\}^{U},\{B,C\}^{U}\}$, $\{\{A\}^{U},\{B,C\}^{L}\}$, $\{\{A\}^{L},\{B,C\}^{U}\}$, $\{\{A\}^{L},\{B,C\}^{L}\}$.
+There are $2^{J(2)}=2^{2}=4$ possibilities to label the cylinders of $P_{2}$ as upper or lower cylinders: $\{\{A\}^{U},\{B,C\}^{U}\}$, $\{\{A\}^{U},\{B,C\}^{L}\}$, $\{\{A\}^{L},\{B,C\}^{U}\}$, $\{\{A\}^{L},\{B,C\}^{L}\}$.
 Example 5.
@@
 That is, if the leveled graph has only three levels and there is only one fragment on level $b=3$, for $c=1$ the algorithm steps through the following partitions:
-$\mathcal{P}_{1}$ =$=\{(A_{3,u}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{2}$ =$=\{(A_{3,u}^{1}/A_{3,l}^{1})^{U}\},$ $\mathcal{P}_{3}$ =$=\{(A_{3,l}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{4}$ =$=\{(A_{3,l}^{1}/A_{3,l}^{1})^{U}\},$
-$\mathcal{P}_{5}$ =$=\{(A_{3,u}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{6}$ =$=\{(A_{3,u}^{1}/A_{3,l}^{1})^{L}\},$ $\mathcal{P}_{7}$ =$=\{(A_{3,l}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{8}$ =$=\{(A_{3,l}^{1}/A_{3,l}^{1})^{L}\}.$
+$\mathcal{P}_{1}$ $=\{(A_{3,u}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{2}$ $=\{(A_{3,u}^{1}/A_{3,l}^{1})^{U}\},$ $\mathcal{P}_{3}$ $=\{(A_{3,l}^{1}/A_{3,u}^{1})^{U}\},$
+$\mathcal{P}_{4}$ $=\{(A_{3,l}^{1}/A_{3,l}^{1})^{U}\},$ $\mathcal{P}_{5}$ $=\{(A_{3,u}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{6}$ $=\{(A_{3,u}^{1}/A_{3,l}^{1})^{L}\},$
+$\mathcal{P}_{7}$ $=\{(A_{3,l}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{8}$ $=\{(A_{3,l}^{1}/A_{3,l}^{1})^{L}\}.$
 The following theorem is a consequence of the results of the previous sections.
@@
 cylinders) to build the surface.
-Note that an $n$-punctured sphere can be obtained by gluing $n-1$ cylinders together.
+Note that an $n$-punctured sphere can be obtained by gluing n1n-1 cylinders together.
 Algorithm 2.
 Let $\mathcal{G}$ be a leveled spatial graph with spine $C$.
-Let $e$ be any edge of $C$ and choose a spanning tree $T$ of $\mathcal{G}$ which contains $C-e$.
-The choice of $T-(C-e)$ is independent of the choice of $e$.
+Let $e$ be any edge of $C$ and choose a spanning tree $T$ of $\mathcal{G}$ which contains CeC-e.
+The choice of T(Ce)T-(C-e) is independent of the choice of $e$.
 Embed $T\cup e$ in a sphere $\mathcal{S}_{0}$, such that $C$ is the equator of $\mathcal{S}_{0}$.
 This embedding is cellular, because $T\cup e$ is a connected trivial spatial graph.
-Contracting all edges in $T-C$ gives a hamiltonian leveled spatial graph $\mathcal{G}^{\prime}$ to which algorithm 1 can be applied.
+Contracting all edges in TCT-C gives a hamiltonian leveled spatial graph $\mathcal{G}^{\prime}$ to which Algorithm 1 can be applied.
 If the algorithm terminates successfully, a cellular embedding $(\mathcal{G}^{\prime},\mathcal{S}^{\prime})$ is obtained.
@@
 Note that the underlying abstract graph of a leveled embedding consists of at most one non-planar 2-connected component, by definition.
-For example, the graph $K_{5}\_K_{5}$, obtained by joining two copies of $K_{5}$ by an edge (see fig.
-11), does not admit a leveled embedding, because it contains two non-planar 2-connected components.
+For example, the graph $K_{5}\_K_{5}$, obtained by joining two copies of $K_{5}$ by an edge (see Figure 13), does not admit a leveled embedding, because it contains two non-planar 2-connected components.
 However, if multiple leveled embeddings are attached to each other via trivial spatial graphs, the algorithms presented before can still be applied.
@@
 For each vertex $v_{1}^{i}$ adjacent to $v_{2}^{j}$, identify some vertices of $H_{j}$ with some vertices of the block $\mathcal{G}_{i}$ such that identified vertices belong to exactly one planar graph and one block.
-(See fig.
-10)
-Figure 10: A multi-leveled embedding with three blocks and two connected planar graphs. On the top left, the tree associated to the embedding.
+(See Figure 11 and Figure 12)
+Figure 11: A multi-leveled embedding with three blocks and two connected planar graphs.
+Figure 12: The tree associated to the embedding of Figure 11.
 Figure 13: A multi-leveled embedding of the graph $K_{5}\_K_{5}$ with two blocks, each isomorphic to $K_{5}$, and one planar graph, isomorphic to $K_{2}$.
@@
 As an example, consider the spatial graph $\mathcal{G}$ described in Example 1.
-Although for the given spine the algorithm does not find a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds, it is possible to choose a non-hamiltonian spine for $\mathcal{G}$, shown in fig.
-12, for which algorithm 2 terminates successfully.
-Furthermore, a different hamiltonian spine for $\mathcal{G}$ can be chosen, shown in fig.
-13, which yields a leveled embedding with five levels.
+Although for the given spine the algorithm does not find a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds, it is possible to choose a non-hamiltonian spine for $\mathcal{G}$, shown in Figure 14, for which Algorithm 2 terminates successfully.
+Furthermore, a different hamiltonian spine for $\mathcal{G}$ can be chosen, shown in Figure 15, which yields a leveled embedding with five levels.
 Corollary 2 guarantees the existence of a surface where $\mathcal{G}$ cellular embeds, and indeed Algorithm 1 terminates successfully if applied to this leveled embedding.
-Note that it follows from this example that all hamiltonian leveled spatial graphs with a spine list of the form $(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},\dots,n^{1},n^{2},n^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},\dots,n^{3},n^{2},n^{1})$ for $n\in\mathbb{N}$ can be cellular embedded by applying algorithm 1 to a different choice of the spine, analogously to the example.
+Note that it follows from this example that all hamiltonian leveled spatial graphs with a spine list of the form $(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},\dots,n^{1},n^{2},n^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},\dots,n^{3},n^{2},n^{1})$, for $n\in\mathbb{N}$, can be cellular embedded by applying Algorithm 1 to a different choice of the spine, analogously to the example.
 Figure 14: Left: the new spine chosen for $\mathcal{G}$ in red. Right: the leveled embedding with the new spine.
@@
 Indeed, the trefoil knot with an unknotting tunnel is an example of a spatial graph that does not admit a leveled embedding, because for every unknotted cycle chosen as spine, the fragment will cross the spine non-trivially.
-It does though have a cellular embedding in a torus, as shown in fig.
-14.
+It does though have a cellular embedding in a torus, as shown in Figure 16.
 Figure 16: A trefoil knot with unknotting tunnel cellular embeds in a torus but does not admit a leveled embedding.
-In [1], it is shown that any spatial graph $\mathcal{G}$ can be embedded in a 3-book, where the vertices lie on the back of the 3-book and edges may be contained in several leaves (i.e.
+In Atn (68), it is shown that any spatial graph $\mathcal{G}$ can be embedded in a 3-book, where the vertices lie on the back of the 3-book and edges may be contained in several leaves (i.e.
 the pages of the 3-book).
@@
 If a spatial graph $\mathcal{G}$ has a leveled embedding, then a cycle can be found that is the spine of a leveled embedding of $\mathcal{G}$ for which Algorithm 2 ends successfully.
-A similar conjecture can be stated for hamiltonian leveled spatial graphs. The situation shown in fig. 13 is representative for all examples of hamiltonian leveled spatial graphs that we studied.
+A similar conjecture can be stated for hamiltonian leveled spatial graphs. The situation shown in Figure 15 is representative for all examples of hamiltonian leveled spatial graphs that we studied.
 Conjecture 3.
@@
 A circular embedding is a cellular embedding $(\mathcal{G},\mathcal{S})$ where the boundaries of the faces of $\mathcal{S}\setminus\mathcal{G}$ are cycles of $\mathcal{G}$.
-A famous open conjecture in graph theory states that every 2-connected graph has a circular embedding [6].
+A famous open conjecture in graph theory states that every 2-connected graph has a circular embedding Jae (85).
 In case of a 2-connected hamiltonian leveled spatial graph $\mathcal{G}$ for which Algorithm 1 ends successfully, it is possible to determine whether the resulting cellular embedding is circular by applying 1 to all faces of the cellular embedding: if every edge of $\mathcal{G}$ appears in the boundary of exactly two faces, then the embedding is circular.